### v1

# Hier-SLAM++: Neuro-Symbolic Semantic SLAM with a Hierarchically Categorical Gaussian Splatting

Abstract
We propose Hier-SLAM++, a comprehensive Neuro-Symbolic semantic 3D Gaussian Splatting SLAM method with both RGB-D and monocular input featuring an advanced hierarchical categorical representation, which enables accurate pose estimation as well as global 3D semantic mapping. The parameter usage in semantic SLAM systems increases significantly with the growing complexity of the environment, making scene understanding particularly challenging and costly. To address this problem, we introduce a novel and general hierarchical representation that encodes both semantic and geometric information in a compact form into 3D Gaussian Splatting, leveraging the capabilities of large language models (LLMs) as well as the 3D generative model. By utilizing the proposed hierarchical tree structure, semantic information is symbolically represented and learned in an end-to-end manner. We further introduce a novel semantic loss designed to optimize hierarchical semantic information through both inter-level and cross-level optimization. Additionally, we propose an improved SLAM system to support both RGB-D and monocular inputs using a feed-forward model. To the best of our knowledge, this is the first semantic monocular Gaussian Splatting SLAM system, significantly reducing sensor requirements for 3D semantic understanding and broadening the applicability of semantic Gaussian SLAM system. We conduct experiments on both synthetic and real-world datasets, demonstrating superior or on-par performance with state-of-the-art NeRF-based and Gaussian-based SLAM systems, while significantly reducing storage and training time requirements.
Index Terms:
Semantic SLAM, hierarchical category, gaussian splatting, RGB-D, monocular.
I Introduction
Visual Simultaneous Localization and Mapping (SLAM) is a key technology for ego-motion estimation and scene perception, widely employed in applications across various domains, such as autonomous drones [1], self-driving vehicles [2], and interactive experiences in Augmented Reality (AR) and Virtual Reality (VR) [3]. Semantic information, which provides high-level knowledge about an environment, is indispensable for comprehensive scene understanding. Thus, it plays a crucial role in enabling intelligent robots to perform various complex tasks. In recent decades, advances in image segmentation and enhanced map representations have accelerated progress in semantic visual SLAM [4, 5, 6, 7, 8, 9, 10, 11, 7, 12].
Figure 1: (a). The hierarchical tree is generated by integrating geometric information and semantic messages, utilizing 3D generative models and Large Language Models (LLMs). (b). The global 3D Gaussian map generated by Hier-SLAM++ with learned semantic labels is shown on the left. The established hierarchical tree of the semantic information is organized on the right. Based on the established tree, the hierarchical symbolic representation for semantic information is shown at the bottom of the block, which compresses semantic data, reducing both memory usage and training time of the semantic SLAM. (c). The rendered semantic map at different levels shows a coarse-to-fine understanding, beneficial for real-world scenarios with shifting perspectives from distant to close.
Recently, 3D Gaussian Splatting has gained attention as a leading 3D world representation method [13, 14, 15], thanks to its fast rendering and optimization efficiency enabled by the highly parallelized rasterization of 3D primitives. Specifically, 3D Gaussian Splatting models the continuous distributions of geometric parameters through Gaussian distributions. This approach not only improves the quality of 3D representation but also supports efficient optimization, making it particularly well-suited for SLAM tasks. While promising, current SLAM systems based on 3D Gaussian Splatting [16, 17, 18, 19, 20] primarily focus on geometric reconstruction. However, the lack of semantic integration limits its potential for complex tasks such as visual navigation, planning, and autonomous driving. Addressing this gap by incorporating semantic understanding without sacrificing efficiency remains a key challenge.
A naive solution would be to augment 3D primitives with discrete semantic labels and parameterize their distribution using a categorical distribution, i.e., a flat Softmax-based representation. However, 3D Gaussian Splatting is already a storage-intensive approach [21, 22], requiring a large number of 3D primitives with multiple parameters to achieve realistic rendering. Adding semantic distribution parameters would further increase storage requirements and processing time, with grow linear in the number of semantic classes, making it particularly impractical for complex scene understanding. Furthermore, given a fixed number of observations (video frames), increasing the number of optimizable parameters leads to the curse of dimensionality, resulting in degraded optimization performance in high-dimensional spaces, such as when handling a large number of semantic classes. Overall, this added overhead for semantic learning is especially problematic for robotics applications, where real-time performance and limited resources are critical. Recent studies have explored alternative methods to mitigate these challenges. For instance, [23] directly learns a 3-channel RGB visualization to avoid semantic labels learning. [24] adopts a flat semantic representation and relies on a pre-trained foundation model to generate an embedding feature map rather than an explicit 3D semantic label map.
Unlike flat representations, semantic information naturally organizes into a hierarchical structure of classes, as shown in Fig. 1. The attributes (such as structures) and properties (such as surfaces) of the semantic classes are extracted as symbolic nodes to generate a hierarchical tree, where each semantic class is represented as a hierarchical expression based on a root-to-leaf path. This hierarchy enable efficient coding of extensive information with a relatively small number of symbolic nodes to learn, resulting in a compact representation. For example, a binary tree with a depth of $10$ can represent $2^{10}$ classes, allowing for the encoding of 1,024 classes using only 20 codes (i.e., $2\times 10$ through 2-dimensional Softmax coding at each level), enabling a reduction of up to $O(\log N)$. Building upon this concept, Hier-SLAM [25] was first introduced as an RGB-D Gaussian Splatting SLAM that employs a hierarchical tree to represent semantic information in an unknown environment, achieving strong performance with improved storage efficiency and operational speed.
Hier-SLAM [25] have demonstrated the advantages and potential of incorporating hierarchical categorization for a semantic RGB-D SLAM system. However, it is limited by a simplistic hierarchical tree representation and lacks robust semantic optimization. To address these limitations and further improve its effectiveness, we propose Hier-SLAM++, a comprehensive neuro-symbolic semantic Gaussian Splatting SLAM framework that incorporates an advanced hierarchical representation for semantic information. Different from Hier-SLAM, we incorporate both geometric (i.e., shape and size details) and semantic information into the hierarchical symbolic tree expression using both Large Language Models (LLMs) and 3D Generative models. Specifically, we utilize a text-to-3D generative model [26] to extract shape information for each class and organize it into hierarchical levels. Additionally, we leverage LLMs to include both semantic and size information in the tree structure, effectively addressing the lack of size information in 3D generative models. With the established tree, semantic information in the environment is represented as symbolic nodes along root-to-leaf paths, to be learned in an end-to-end manner during the SLAM operation. To enhance semantic understanding, we introduce an improved hierarchical loss function that incorporates both inter-level and cross-level optimizations. This coarse-to-fine hierarchical approach based on the established hierarchical representation is beneficial for real-world applications, especially those involving observations from distant to nearby views. Furthermore, our SLAM system is extended to support both RGB-D and monocular inputs. We utilize the geometric output of the 3D feed-forward model [27] as a geometric prior, removing the depth dependency of existing SLAM methods. This enables Hier-SLAM++ to perform semantic SLAM without relying on dedicated depth sensors, broadening the range of applicable scenarios. Finally, we enhance and refine the 3D Gaussian Splatting SLAM to improve both performance and efficiency. We conduct experiments on multiple synthetic and real-world datasets. The results demonstrate that our SLAM system outperforms or is on par with state-of-the-art NeRF-based and Gaussian-based SLAM methods. Additionally, our approach effectively reduces storage and training time.
The technical contributions of our work include:
1) We introduce an advanced hierarchical tree representation that integrates semantic, shape, and size information, leveraging the capabilities of both LLMs and 3D generative models. This tree-based coding efficiently compresses semantic information while maintaining its inherent hierarchical structure, resulting in reduced memory usage and shorter training times.
2) We present a novel optimization approach for the proposed semantic hierarchical representation, integrating both inter-level and cross-level losses to ensure thorough optimization across all levels of the hierarchical semantic information.
3) We propose an improved semantic SLAM system capable of supporting both RGB-D and monocular inputs by leveraging a 3D feed-forward model. To the best of our knowledge, Hier-SLAM++ is the first Gaussian Splatting based semantic SLAM system to support monocular input, significantly reducing sensor requirements and broadening its application scope.
4) We conduct extensive experiments on both synthetic and real-world datasets, comprehensively demonstrating the effectiveness of our method and its improvements.
II Related Work
II-A Visual SLAM system
Visual Simultaneous Localization and Mapping (SLAM) using visual sensors is a long-standing and active research topic that focuses on the simultaneous estimation of ego-motion and the reconstruction of a global 3D map. Traditional SLAM methods [28, 29, 30, 31], developed over many years, have demonstrated both high accuracy and robustness, even in challenging environments. This makes them an indispensable frontend foundation for various real-world applications, such as autonomous driving and drones. In recent years, several new 3D representations have emerged in the 3D computer vision fields. With the advent of neural implicit representations, numerous works [32, 33, 34, 35] have demonstrated the effectiveness of NeRF-based SLAM approaches. These new representations introduce novel capabilities to the SLAM community, including enhanced rendering capability. 3D Gaussian Splatting is the latest advancement in the 3D vision field, offering significantly faster rendering speed and strong performance compared with neural implicit representations. Recent works [16, 17, 20, 19] have demonstrated new SLAM capabilities by integrating 3D Gaussian Splatting techniques, further pushing the boundaries of real-time scene understanding and reconstruction.
II-B Neural implicit visual SLAM
Neural Radiance Fields (NeRF) [36] is a neural rendering technique that represents 3D scenes using a fully connected neural network. iMap [32] proposes a real-time keyframe-based RGB-D SLAM method that uses a multilayer perceptron (MLP) as the sole scene representation. Nice-SLAM [33] introduces a dense SLAM system that incorporates multi-level local information by employing a hierarchical scene representation. Nicer-SLAM [34] extends this approach by proposing a monocular SLAM system with a hierarchical neural implicit map representation, eliminating the dependency on depth information. Integrating SLAM with semantic understanding, many works [7, 11, 12] have explored neural implicit representations for semantic mapping and localization tasks. DNS-SLAM [7] incorporates 2D semantic priors with a coarse-to-fine geometry representation to embed semantic information into the map. SNI-SLAM [11] integrates geometry, appearance, and semantic features into a shared feature space to enhance the robustness of the SLAM method. However, these approaches are constrained by the slow convergence of neural implicit map representations, which often leads to performance degradation when coupled with semantic objectives [37, 38]. In contrast, Gaussian Splatting provides significant advantages, including fast rendering performance and high-density reconstruction quality, making it a promising alternative for map representation.
II-C 3D Gaussian Splatting SLAM
3D Gaussian Splatting has recently gained attention as a promising 3D representation due to its fast rendering performance and high-density reconstruction quality. Leveraging this 3D representation, SplaTAM [16] uses silhouette guidance for pose estimation and map reconstruction in RGB-D SLAM systems, while MonoGS [17] supports both monocular and RGB-D SLAM. Photo-SLAM [19] utilizes 3D points generated by ORB-SLAM as position parameters for Gaussian Splatting primitives within the SLAM system. These studies highlight the strong potential of 3D Gaussian Splatting in SLAM tasks across various scenarios [16, 17, 18, 19, 20]. However, integrating semantic understanding into SLAM presents a significant challenge, as it requires jointly solving three high-dimensional optimization problems with varying value ranges and convergence behaviors, while also significantly increasing storage requirements. Focus on the integration of semantic information into the Gaussian Splatting SLAM, SGS-SLAM [23] added RGB 3-channels to create a semantic visualization map to avoid true semantic labels learning. SemGauss-SLAM [24] uses a flat semantic representation supervised by a large pre-trained foundation model. However, these methods overlook the natural hierarchical structure of the real world. Additionally, relying on large foundation models increases the complexity of the neural network and its computational demands, with performance largely dependent on the embeddings from these pre-trained models. Eliminating the dependency on foundation models, this work proposes a simple yet effective hierarchical representation for semantic understanding in the SLAM system which largely reduces the training storage and time.
III Method
Figure 2: Left: Overview of the Hier-SLAM++ pipeline. The global 3D Gaussian map is initialized using the first frame of the video stream input. The system then alternates between the Tracking and Mapping steps as new frames are processed (see Section III-D). In the RGB-D setting, depth is directly obtained from sensor input, whereas in the monocular setting the 3D feed-forward method (DUSt3R) is used to generate a geometric prior for depth estimation (see Section III-C). Top Right: Hierarchical representation of semantic information. The Tree Generation process leverages the capabilities of both LLMs and 3D Generative Models in a top-to-bottom manner. This hierarchical tree is used to establish a symbolic coding for each Gaussian primitive (see Section III-A). Additionally, we introduce a novel loss function that combines Inter-level Loss $L_{\text{Inter}}$ and Cross-level Loss $L_{\text{Cross}}$ to optimize the hierarchical semantic representation (see Section III-B). Bottom Right: An example of hierarchical semantic rendering.
The entire pipeline of our method is illustrated in Fig. 2. Within Hier-SLAM++, we model the semantic space as a hierarchical tree, where semantic information is represented as a root-to-leaf hierarchical symbolic structure. This representation not only comprehensively captures the hierarchical attributes of semantic information, but also effectively compresses storage requirements, while enhancing semantic understanding with valuable scaling-up capability. To construct an effective tree, we leverage both LLMs and 3D Generative Models to integrate semantic and geometric information. Next, each 3D Gaussian primitive is augmented with a hierarchical semantic embedding code, which can be learned in an end-to-end manner through the proposed Inter-level and Cross-level semantic losses. Furthermore, our method is extended to support monocular input by incorporating geometric priors extracted from the feed-forward method, thereby eliminating the dependency on depth information. The following sections provide detailed descriptions for each part of our Hier-SLAM++.
III-A Hierarchical representation
We propose hierarchical symbolic representation for semantic information to provide improved efficiency, compression, and compactness for semantic SLAM, making it well-suited for multi-parameter joint optimization and enabling scalability to complex scenes. In this section, we provide a detailed explanation of the hierarchical tree parametrization and its generation process.
III-A1 Tree Parametrization
The hierarchical tree $\bm{G}$ adopted in Hier-SLAM++ is represented as the set of vertices and edges, $\bm{G}=(\bm{V},\bm{E})$. The vertices set $\bm{V}=\cup_{l=0}^{L}\left\{\bm{v}_{l}\right\}$ comes from the classes from all tree levels, where $\left\{\bm{v}_{l}\right\}$ represents the set of nodes at the $l$-th level of the tree, and we use $L$ to represent depth of the whole tree. The edge set $\bm{E}=\cup_{m=0}^{L-1}\left\{\bm{e}_{m}\right\}$ captures the subordinate relationships, reflecting semantic attribution, size message, and geometric knowledge. For the $i$-th semantic class $g^{i}$, which is treated as a single leaf node in the tree view, its hierarchical expression is:
$g^{i}=\{v_{l}^{i},e_{m}^{i}\mid l=0,1,...,L;\;m=0,1,...,L-1\},$ (1)
which corresponds to the root-to-leaf path: $g^{i}=v_{0}^{i}\xrightarrow{e_{0}}v_{1}^{i}\xrightarrow{e_{1}}\cdots\xrightarrow{e_{L-2}}v_{L-1}^{i}\xrightarrow{e_{L-1}}v_{L}^{i}$, as illustrated in Fig. 3. Take a semantic class $g=\texttt{Bed}$ as an example, a 5-level tree symbolic coding can be represented as: $\left\{v_{0}:\texttt{LargeItems}\right\}\rightarrow\left\{v_{1}:\texttt{Furnishings}\right\}\rightarrow\left\{v_{2}:\texttt{BedroomFurniture}\right\}\rightarrow\left\{v_{3}:\texttt{Rectangular}\right\}\rightarrow\left\{v_{4}:\texttt{Bed}\right\}$. The hierarchical attributes and properties of class $g^{i}$ are represented by the symbolic node information $v_{l}^{i}$, while its relationships such as ‘include’ and ‘possessing’ are represented by the edge information $e_{m}^{i}:\rightarrow$. In this way, every semantic class can be coded in a progressive, symbolic hierarchical manner, incorporating both semantic and geometric perspectives. Moreover, the standard flat representation can be seen as a single-level tree coding from the hierarchical viewpoint. Herein, we shall omit the superscripts (class indexes) of variables for clarity and compact notation.
III-A2 Tree Generation with LLMs and 3D Generative Models
To represent semantic information both accurately and comprehensively, an effective tree structure should incorporate information from both semantic and geometric perspectives, as these are the two crucial types of information in an unknown environment, which are essential for achieving a comprehensive understanding of the global 3D environment. For capturing semantic meanings, such as functional relationships and summary description labels, we use Large Language Models (LLMs), specifically GPT-4 Turbo [39], known for its strong commonsense reasoning and language understanding capabilities. On the other hand, to capture general geometric information without dataset limitations, we employ 3D Generative Models, specifically text-to-3D models, as they have the capability to generate generalized and unified 3D objects with accurate shape information based on text prompts after training. We adopt the text-to-3D approach MeshGPT [26], which learns 3D symbolic embeddings with the local mesh geometry and topology of 3D assets, making it well-suited for our purpose. However, the text-to-3D method only generates 3D objects within a unified coordinate system, omitting size information. To address this issue, we complement the geometric shape information with the size attributes derived from LLMs, to capture the comprehensive geometric information for each semantic class. The leverage of both LLMs and 3D Generative Models enables the creation of an elaborate and well-structured hierarchical tree representation, where each level reflects a distinct aspect of a class’s attributes, enhancing the depth and utility of the global 3D understanding. The following part provides a detailed explanation of the tree generation process.
As illustrated in Fig. 3, the tree generation process follows a top-to-bottom manner. At the upper level, the initial grouping divides all classes into several subgroups based on their typical physical sizes in scenes, obtained through prompt queries to the LLM:
$\{v,e\}_{0}=\mathcal{L}^{\text{size}}\{g\}$ (2)
where $\mathcal{L}^{\text{size}}$ represents the LLMs usage for the size message, and the subscript in $\{v,e\}_{0}$ represents the generation at the initial levels. The clustering results categorize classes into groups such as small items, medium items, and large items. We provide details of the specific prompt used for the LLM in Appendix 1.
Next, the classes within each group are used as input prompts to LLMs for further clustering based on their semantic relationships, particularly their functional properties:
$\{v,e\}_{1}=\mathcal{L}^{\text{func}}\{g\}$ (3)
where $\mathcal{L}^{\text{func}}$ denotes the LLMs clustering based on functional relationships, and the subscript of $\{v,e\}_{1}$ stands for the following levels generation. We demonstrate the detail prompt used for the LLM in Appendix 1.
After multiple levels generation using LLMs, 3D Generative Models are employed to cluster classes based on their geometric shapes. Specifically, we adopt MeshGPT [26] to capture the geometric information for each semantic class. MeshGPT [26] is a text-to-3D model that encodes object shape and internal topology relationships into latent quantized embeddings, forming a symbolic 3D representation in the shape space. Given a text prompt (i.e., the class name) as input to the pre-trained MeshGPT, we extract the latent quantized embeddings to represent the shape information for the class:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{z}=\mathcal{E}\{g\}\in\mathbb{N}^{F\times 2}}}$ (4)
where $g$ is the class name provided as input. $\mathcal{E}$ represents the pre-trained MeshGPT encoder, and $\bm{z}$ is the quantized geometric embedding extracted from the pre-trained encoder. For each object’s embedding, $F$ is the number of triangle faces used in the mesh representation. Since the number of triangle faces varies across different semantic objects, i.e., different values of $F$, we first compute the average shape embedding $\bm{z}_{\text{avg}}\in\mathbb{R}^{1\times 2}$ for each object. The averaged embeddings from all classes are then concatenated and subsequently normalized to a standard distribution (zero mean and unit variance) to ensure consistency across classes:
$\tilde{\bm{z}}=\mathcal{N}\{\mathcal{C}(\bm{z}_{\text{avg}})\}$ (5)
where $\mathcal{N}$ and $\mathcal{C}$ denote the normalization and concatenation operations, respectively. Next, the normalized embeddings $\tilde{\bm{z}}$ are clustered using K-means++, denoted by the operator $\mathcal{K}$. To address the lack of descriptive labels for each clustering result, we further employ LLMs to summarize and refine the clustering outcomes:
$\{v,e\}_{2}=\mathcal{L}^{\text{sum}}\{\mathcal{K}(\tilde{\bm{z}})\}$ (6)
where $\mathcal{L}^{\text{sum}}$ represents the use of LLMs for summarization. The generated descriptive labels, such as ”box”, ”flat”, ”soft”, summarize the common geometric attributes within each grouping result, further enhancing the interpretability of the hierarchical tree. The details are illustrated in Appendix 1.
Whenever LLMs are employed, we ensure balanced, meaningful, and descriptive outputs through carefully designed prompt inputs, as detailed in Appendix 1. Furthermore, an additional LLM validator [25] is utilized after each generation level to verify clustering completeness and accuracy, ensuring that no classes are missing or incorrectly included (see Appendix 1 for detail). Finally, a global tree evaluation is conducted to detect and resolve any duplicates or omissions across tree levels, ensuring a fully balanced and complete hierarchical tree structure. We use manual inspection as the final verification to ensure the correctness and rationality of the holistically generated tree.
Figure 3: Visualization of the hierarchical tree structure and semantic embedding. For tree representation, each semantic class is expressed hierarchically as $g=\{v_{l},e_{m}\}$, where edges $e_{m}$ are depicted as lines connecting tree nodes. The nodes within the same level are illustrated via the same color. For tree generation, we utilize both LLMs and 3D Generative Models to extract and group messages. For hierarchical semantic embedding, we represent semantic information using a root-to-leaf symbolic path. We propose two types of representations: the one-hot representation and the binary representation, which can reduce the original dimensionality by up to $O(\log^{2}N)$.
III-B Hierarchical optimization
Based on the generated hierarchical tree structure, the semantic information is represented as compact hierarchical symbolic nodes, and further integrated into the 3D Gaussian primitives, learned in an end-to-end manner with the input frame stream. To ensure accurate semantic understanding, we introduce a semantic hierarchical loss, which consists of Inter-level and Cross-level losses. These losses are incorporated into the optimization process during the online SLAM operation. The following sections provide a detailed explanation of the proposed tree encoding and the loss calculation.
III-B1 Tree Encoding
As introduced in Eq. (1), the semantic class in Hier-SLAM++ is represented hierarchically in a root-to-leaf symbolic manner based on the generated tree. This effectively compacts semantic information, facilitates SLAM joint optimization, and provides valuable scalability for semantic understanding in unknown environments. Specifically, we propose two types of hierarchical semantic embedding approaches, both of which compact the original semantic information.
One-hot representation: As demonstrated in Fig. 3, the hierarchical semantic embedding $\bm{h}$ is composed of the one-hot embeddings across all levels:
$\bm{h}=\mathcal{C}(\bm{h}_{l})\in\mathbb{B}^{N},\quad\bm{h}_{l}\in\mathbb{B}^{n}$ (7)
where $l$ represents the $l$-th level of the tree, and $\mathcal{C}$ denotes the concatenation operation. For each level’s representation $\bm{h}_{l}$, we use $n$-dimensional one-hot codes (in colored boxes) to encode the symbolic nodes (in colored circles) from the tree of this level, as shown in Fig. 3. The dimension $n$ corresponds to the maximum number of nodes at the $l$-th level. Using the one-hot representation, the overall dimension of the semantic embedding is the sum of the dimensions across all levels, given by $N=\sum n$. This enable a maximum reduction of up to $O(\log N)$ compared to the original dimension $N_{\text{original}}$ in the flat representation. This hierarchical representation effectively compacts the semantic embedding, reducing storage requirements and facilitating joint optimization.
Binary representation: We propose a further compact version of our hierarchical semantic representation in this section. The binary representation of the hierarchical semantic embedding $\bm{b}$ consists of the binary embeddings $\bm{b}_{l}$ across all tree levels:
$\bm{b}=\mathcal{C}(\bm{b}_{l})\in\mathbb{B}^{K},\quad\bm{b}_{l}\in\mathbb{B}^{k}$ (8)
where $\bm{b}_{l}$ is the binary representation converted from the one-hot semantic embedding $\bm{h}_{l}$ at each level: $\bm{h}_{l}\rightarrow\bm{b}_{l}$, as depicted in the last row example in Fig. 3. For instance, if the one-hot embedding has a dimension of $n=8$, this corresponds to a maximum of $8$ unique codes at this level (covering values from 00 to $7$). After binarization, the dimension of the compact version is reduced to $k=3$, since $82^{3}=$ 8. By leveraging binary representation, the hierarchical coding dimension at each level is reduced to $k=\log n$, compared to the one-hot encoding with dimension $n$. Consequently, the total semantic encoding dimension across all levels is given by $K=\sum k$, resulting in a more compact representation.
III-B2 Loss Calculation
To fully optimize the hierarchical semantic message both within each level and across different levels of the tree, the proposed hierarchical loss consists of Inter-level loss $L_{\text{Inter}}$ and Cross-level loss $L_{\text{Cross}}$:
$L_{\text{{Semantic}}}=\omega_{1}L_{\text{{Inter}}}+\omega_{2}L_{\text{{Cross}}}$ (9)
where $\omega_{1}$ and $\omega_{2}$ are balancing coefficients for each loss term. Specifically, the Inter-level loss $L_{\text{Inter}}$ is employed within each level’s embedding $\bm{h}^{l}$ to ensure the optimization of semantic information within each level:
$L_{\text{{Inter}}}=\sum_{l=0}^{L}L_{\text{{ce}}}(\mathcal{S}(\bm{h}^{l}),\mathcal{P}^{l})$ (10)
where $L_{\text{{ce}}}$ indicates the cross-entropy semantic loss, and $\mathcal{P}^{l}$ denotes the semantic ground truth at level $l$. $\mathcal{S}$ stands for the Softmax operation, converting embeddings into probabilities. For our proposed compact version with binary representation $\bm{b}^{l}$, we replace the original cross-entropy loss with binary cross-entropy loss $L_{\text{bce}}$ for the binary Inter-level loss calculation:
$L_{\text{Inter}}^{\text{bin}}=\sum_{l=0}^{L}L_{\text{{bce}}}(\mathcal{S}(\bm{b}^{l}),\mathcal{P}^{l})$ (11)
To enforce accurate understanding across tree levels, we introduce the Cross-level loss, which is computed using a semantic decoder based on the holistic hierarchical codes $\bm{h}$. Specifically, we first perform a 2D convolution operation on the semantic embedding, following an activation layer (i.e., ReLU). Then, a second 2D convolution layer is applied to map the hidden embedding into the flat coding, followed by a softmax operation to convert the embeddings into probabilities. The overall Cross-level loss $L_{\text{Cross}}$ is defined as:
$L_{\text{Cross}}=L_{\text{ce}}\big{(}\mathcal{S}(\text{decoder}(\bm{h})),\mathcal{P}\big{)},$ (12)
where $\mathcal{P}$ represents the semantic ground truth, and the definition of $\text{decoder}(\bm{h})$ is:
$\text{decoder}(\bm{h})=\text{Conv}\big{(}\text{ReLU}(\text{Conv}(\bm{h}))\big{)}$ (13)
III-C Monocular setting with geometric priority
Accurate depth information provides valuable geometric cues for global map reconstruction and pose estimation in a SLAM system. In an RGB-D setting, depth signals are directly obtained from sensors. However, in a monocular setting without depth information, SLAM [28, 29, 30, 31] typically relies on triangulation to initialize the global mapping. Yet, due to the inherent coupling between pose estimation and mapping, this approach often suffers from relatively lower accuracy compared with RGB-D based methods. In our Hier-SLAM++, we employ a feed-forward method as a geometric prior. Specifically, we leverage DUSt3R [27] to provide depth geometric priors for monocular SLAM. DUSt3R is able to generate both a 3D point map and pose estimations from several sparse-view image inputs, which can serve as the geometric prior for the monocular SLAM. The following section provides details on the monocular setting in our Hier-SLAM++.
III-C1 Geometric prior acquisition
DUSt3R [27] supports sparse view inputs with a minimum of two views. Larger view inputs enhance reconstruction accuracy by providing more information, yet at the cost of increased computational demands. To balance reconstruction accuracy and computational efficiency, we sample every $S$ frames with a skip of $\delta$ frames from the input video to form an image set $\{I_{i}\}$.
$\{I_{i}\;|\;i=i_{0}+\delta\cdot(s-1),\;s=1,\dots,S\}$ (14)
Each image set is then fed into DUSt3R to generate geometric information for every image frame, specifically the depth maps. Although DUSt3R is also able to provide pose predictions for each input frame, we observe that the accuracy of estimated pose is relatively low—performing even worse than a constant velocity pose prediction in the SLAM system based on our experiments. Therefore, only the estimated geometric information from DUSt3R is leveraged in our method.
III-C2 Geometric prior correction
The scale of the estimation for each image set varies and also differs from the scale of the online global map estimation during SLAM operation, requiring an additional geometric prior correction operator.
Firstly, we calculate the initial scale and shift parameters by comparing the rendering results from the current global map prediction with the depth prior obtained from DUSt3R. Specifically, the depth map $D_{c}$ and the silhouette map $S_{c}$ for current frame $I_{c}$ are rendered based on the global Gaussian map $G$ as well as the pose estimation $\mathbf{T}_{c}$ for the current frame:
$\{D_{c},S_{c}\}=\mathcal{R}(G,\mathbf{T}_{c})$ (15)
where $\mathcal{R}$ represents the rasterization procedure for the 3D Gaussian map. Next, the global scale and shift parameters $\lambda_{c},\tau_{c}$ are calculated by comparing the predicted depth $\hat{D}_{c}$ from DUSt3R with the rendered depth $D_{c}$ using the least squares estimation on the visible regions $M_{c}$ defined by $S_{c}$:
$\{\lambda_{c},\tau_{c}\}=\arg\min\left(\hat{D}_{c}^{\text{align}}-D_{c}\right)^{2}M_{c}$ (16)
where $\hat{D}_{c}^{\text{align}}=\lambda_{c}\cdot\hat{D}_{c}+\tau_{c}$.
Secondly, the estimated parameters are fed into the optimization as initialization and further optimized jointly within the map optimization via the depth supervision loss:
$\{\lambda_{c},\tau_{c},G\}=\arg\min|\hat{D}_{c}^{\text{align}}-D_{c}|M_{c}$ (17)
In this way, the depth prediction can be continuously refined online with increasing observations, providing geometric supervisory signals during SLAM optimization. This helps guide global reconstruction and pose estimation, leading to improved results throughout the SLAM process.
III-D Semantic Gaussian Splatting SLAM
The global semantic 3D Gaussian Splatting SLAM system is illustrated in the left part of Fig. 2. Based on the established semantic 3D Gaussian representation, global 3D reconstruction (Mapping) and pose estimation (Tracking) are alternately performed with the video stream input. For the RGB-D setting, depth information for each image frame is obtained directly from depth sensors. For the monocular setting, we employ a feed-forward method to obtain the geometric prior, as introduced in the previous section. Further details of our Hier-SLAM++ system are provided in the following section.
III-D1 Semantic 3D Gaussian representation
The semantic 3D Gaussian representation is represented as the Gaussian primitives with the hierarchical semantic embedding. Motivated by [16], we adopt the isotropic and view-independent Gaussian integrated with the semantic embeddings. Each semantic Gaussian is composed of color $\bm{c}$, the center position $\bm{\mu}$, the radius $r$, the opacity $o$, and its semantic embedding $\bm{h}$. And the influence of each Gaussian $G(\bm{X})$ [13] according to the standard Gaussian equation is:
$G(\bm{X})=o\>\exp\left(-\frac{||\bm{X}-\bm{\mu}||^{2}}{2r^{2}}\right)$ (18)
where $\bm{X}$ stands for the 3D point.
Following [13], every semantic 3D Gaussian primitive within the current global map is projected to the 2D image space using the tile-based differentiable $\alpha$-compositing rendering. The semantic map $H$ is rendered as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{H=\sum_{i=1}^{n}\bm{h}_{i}\alpha_{i}(\bm{X})T_{i}\quad\text{with}\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{X}))}}$ (19)
The rasterized color image $C$, depth image $D$, and the silhouette image $S$ are defined as follows:
$C=\sum_{i=1}^{n}\bm{c}_{i}\alpha_{i}(\bm{X})T_{i},\;D=\sum_{i=1}^{n}\bm{d}_{i}\alpha_{i}(\bm{X})T_{i},\;S=\sum_{i=1}^{n}\alpha_{i}(\bm{X})T_{i}$ (20)
Different from previous work [16], which uses separate forward and backward Gaussian modules for different parameters, our Hier-SLAM++ system adopt a unified forward and backward modules that processes all parameters, including semantics, color, depth, and silhouette images—significantly improving overall runtime efficiency.
III-D2 Tracking
The tracking step focuses on pose estimation for each frame. For each incoming frame, the constant velocity model is first applied to initialize its pose based on the previous frame’s pose estimation. Following that, a pose optimization is performed with the global Gaussian map $G$ fixed:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathbf{T}_{c}=\arg\min_{\mathbf{T}_{c}}L_{\text{Track}}(G,\mathbf{T}_{c})}}$ (21)
where the tracking loss includes the rendering color and depth losses:
$L_{\text{{Track}}}=M\big{(}w_{1}L_{\text{{Depth}}}(G,\mathbf{T}_{c})+w_{2}L_{\text{{Color}}}(G,\mathbf{T}_{c})\big{)}$ (22)
where $L_{\text{Depth}}$ and $L_{\text{Color}}$ denote the L1-loss for the rendered depth and color information. The rendering follows the tile-based differentiable $\alpha$-compositing procedure, utilizing the current pose estimation along with the current global map estimation. Here, weights $w_{1}$ and $w_{2}$ are introduced to balance the two loss terms, and the optimization is only performed on the silhouette-visible image regions, defined as $M=(S>\delta)$. For the RGB-D setting, depth supervisory signals are directly obtained from depth sensors, whereas for the monocular setting, we leverage the geometric prior derived from the feed-forward method.
III-D3 Mapping
The global map information, including semantic information, is optimized in the mapping procedure while keeping camera poses fixed:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{G=\arg\min_{G}L_{\text{Map}}(G,\mathbf{T}_{i})}}$ (23)
The mapping optimization losses include depth, color, and semantic losses:
$L_{\text{{Map}}}=w_{3}ML_{\text{{Depth}}}(G,\mathbf{T}_{i})+w_{4}L_{\text{{Color}}}^{\prime}(G,\mathbf{T}_{i})+w_{5}L_{\text{{Semantic}}}(G,\mathbf{T}_{i})$ (24)
where $L_{\text{{Semantic}}}$ is the proposed semantic loss introduced in Section III-B, and $L_{\text{{Color}}}^{\prime}$ is the weighted sum of SSIM color loss and L1-Loss. The weights $w_{3}$, $w_{4}$, and $w_{5}$ are used to balance the different loss terms. Similar to the tracking process, depth supervisory signals come from sensors in the RGB-D setting. For the monocular setting, geometric priors derived from the feed-forward method are applied, and the two parameters (scale and shift) are further jointly optimized within the mapping process.
IV Experiments
IV-A Experiment settings
The experiments are conducted on both synthetic and real-world datasets, including ScanNet [40], Replica [41], and TUM-RGBD [42]. Following the evaluation metrics which used in previous SLAM methods [16, 33], we leverage ATE RMSE (cm) to quantify SLAM tracking accuracy. For mapping evaluation, Depth L1 (cm) is used to measure accuracy. To assess image rendering quality, we adopt PSNR (dB), SSIM, and LPIPS as evaluation metrics. In line with previous works [24, 7, 11], we assess semantic rendering performance using mIoU (mean Intersection over Union across all classes) to represent global semantic information. To evaluate computational efficiency, we report the runtime of the proposed SLAM method. For comparison, we benchmark our approach against state-of-the-art dense visual SLAM methods, including both NeRF-based and 3D Gaussian SLAM approaches, to demonstrate its effectiveness. Additionally, we compare with state-of-the-art semantic SLAM methods, covering both NeRF-based and Gaussian-based techniques, to highlight its capability in hierarchical semantic representation and scalability. The experiments are conducted in the Nvidia L40S GPU.
For experimental settings, the semantic embedding of each Gaussian primitive is initialized randomly. The semantic optimization loss weights $\omega_{1}$ and $\omega_{2}$ are set to $1.0$ and $0.0$, respectively, for the first $\eta$ iterations, where $\eta$ is set as $15$. Afterwards, $\omega_{1}$ remains $1.0$, while $\omega_{2}$ is increased to $5.0$. This strategy means that we first apply the Inter-level loss to initialize the hierarchical semantic coding, followed by incorporating the Cross-level loss to refine the whole embedding. For tracking loss, we use $\delta=0.99$, with weights $w_{1}=1.0$ and $w_{2}=0.5$. For mapping, the loss weights are set as follows: $w_{3}=1.0$, $w_{4}=0.5$, and $w_{5}=0.2$.
TABLE I: RGB-D Tracking performance ATE RMSE (cm) on the Replica.
Best results are highlighted as FIRST, SECOND.
Methods Avg. R0 R1 R2 Of0 Of1 Of2 Of3 Of4
iMap [32] 4.15 6.33 3.46 2.65 3.31 1.42 7.17 6.32 2.55
NICE-SLAM [33] 1.07 0.97 1.31 1.07 0.88 1.00 1.06 1.10 1.13
Vox-Fusion [43] 3.09 1.37 4.70 1.47 8.48 2.04 2.58 1.11 2.94
co-SLAM [44] 1.06 0.72 0.85 1.02 0.69 0.56 2.12 1.62 0.87
ESLAM [35] 0.63 0.71 0.70 0.52 0.57 0.55 0.58 0.72 0.63
Point-SLAM [45] 0.52 0.61 0.41 0.37 0.38 0.48 0.54 0.69 0.72
MonoGS-RGBD [17] 0.79 0.47 0.43 0.31 0.70 0.57 0.31 0.31 3.2
SplaTAM [16] 0.36 0.31 0.40 0.29 0.47 0.27 0.29 0.32 0.55
[…]
DNS SLAM [7] 0.45 0.49 0.46 0.38 0.34 0.35 0.39 0.62 0.60
SemGauss-SLAM [24] 0.33 0.26 0.42 0.27 0.34 0.17 0.32 0.36 0.49
SGS-SLAM [23] 0.41 0.46 0.45 0.29 0.46 0.23 0.45 0.42 0.55
Hier-SLAM [25] 0.33 0.21 0.49 0.24 0.29 0.16 0.31 0.37 0.53
[…]
Hier-SLAM++ (binary) 0.31 0.23 0.46 0.23 0.29 0.15 0.27 0.34 0.54
IV-B Camera Tracking Accuracy
Tab. I presents our tracking performance on the Replica dataset [41] in the RGB-D setting, compared with state-of-the-art dense SLAM methods, both with and without semantic information. By accurately learning semantic information, our Hier-SLAM++ outperforms all existing methods in 6 out of 8 sequences. For the remaining two sequences, our method achieves nearly the same performance with sub-millimeter level localization error. The lowest average tracking error clearly demonstrate that our approach surpasses other state-of-the-art methods. We present our RGB-D tracking performance on the real-world ScanNet dataset [40] in Tab. II. Compared to the results on the Replica dataset, all methods exhibit degraded performance due to the noisy and sparse depth sensor input, as well as lower color image quality caused by motion blur. Specifically for semantic understanding, the semantic annotations provided by ScanNet are significantly noisier than those in Replica, containing noticeable noise and imprecise boundaries. All of these issues negatively impact the overall performance of semantic SLAM systems. We evaluate all six sequences and show that our method achieves performance comparable to state-of-the-art approaches [25, 16], demonstrating the robustness of our proposed method under these challenging conditions. The tracking performance on the TUM-RGBD dataset is shown in Tab. III. Since the TUM-RGBD dataset does not contain semantic ground truth, we evaluate our method based on tracking accuracy here. Despite the low-quality RGB and depth inputs, as well as the absence of semantic information, all of which could negatively impact overall performance, our method still achieves similar performance to the state-of-the-art approach [16], with a localization error of less than $0.2$ cm, while operating at a significantly faster runtime. A detailed analysis of the runtime performance will be provided in the following section.
TABLE II: RGB-D Tracking performance ATE RMSE (cm) on the Scannet.
Best results are highlighted as first, second, third,fourth.
Methods Avg. $0$ 0059 0106 0169 0181 0207
NICE-SLAM [33] 10.70 12.00 14.00 7.90 10.90 13.40 6.20
Vox-Fusion [43] 26.90 68.84 24.18 8.41 27.28 23.30 9.41
Point-SLAM [45] 12.19 10.24 7.81 8.65 22.16 14.77 9.54
SplaTAM [16] 11.88 12.83 10.10 17.72 12.08 11.10 7.46
SGS-SLAM [23] 9.87 11.15 9.54 10.43 10.70 11.28 6.11
SemGauss-SLAM [24] 11.87 7.97 8.70 9.78 8.97
Hier-SLAM [25] 11.36 11.45 9.61 17.80 11.93 10.04 7.32
[…]
Hier-SLAM++ (binary) 11.59 13.24 9.60 17.83 11.75 9.83 7.29
TABLE III: RGB-D Tracking performance ATE RMSE (cm) on the TUM-RGBD. Best results are highlighted as first, second, third, fourth.
Methods Avg. fr1/desk fr1/desk2 fr1/room fr2/xyz fr3/off
Kintinuous [46] 4.84 3.70 7.10 7.50 2.90 3.00
ElasticFusion [47] 6.91 2.53 6.83 21.49 1.17 2.52
ORB-SLAM2 [29] 1.98 1.60 2.20 4.70 0.40 1.00
NICE-SLAM [33] 15.87 4.26 4.99 34.49 31.73 3.87
Vox-Fusion [43] 11.31 3.52 6.00 19.53 1.49 26.01
Point-SLAM [45] 8.92 4.34 4.54 30.92 1.31 3.48
SplaTAM [16] 5.48 3.35 6.54 11.13 1.24 5.16
[…]
IV-C Global Mapping Performance
Following previous methods [25, 23, 24], we evaluate the depth error on the Replica dataset [41] to assess overall mapping performance, as it provides the highest-quality depth maps for comparison. With an accurate understanding of semantics, our method effectively learns precise global geometry in unknown environments, benefiting both tracking and mapping in the semantic SLAM system. Our method achieves superior results in 7 out of 8 sequences and maintains a low average depth error across all sequences, with a difference of approximately $0.1$ cm compared to the best performing. These results demonstrate that our approach have strong semantic mapping capability, highlighting our superior mapping performance.
IV-D Rendering Quality
We present the quantitative analysis of the rendering performance of our proposed method in Tab. V. Following previous works, Point-SLAM [45] and NICE-SLAM [33], the evaluation is conducted on the input views. As shown in Tab. V, our method achieves superior image rendering performance compared to state-of-the-art approaches across all 8 sequences, as well as in the overall average evaluation. As discussed in previous sections, incorporating semantic information introduces a significant computational burden to the joint optimization process within the SLAM system. However, our method demonstrates notably improved rendering performance compared to all existing methods, including the non-semantic SLAM approaches. Although the additional computation for semantic learning increases the optimization complexity, the accurate semantic understanding enhances both geometric and appearance learning through joint optimization, effectively improving rendering performance.
The visual comparison of rendered images is presented in Fig. 6. Benefiting from precise semantic information learning and the improved SLAM system, the rendering results exhibit significantly improved detail preservation, such as more accurate and well-defined lamp boundaries in the first and fifth columns, and precise details even in fine structures like shutters in the second column. Moreover, our method produces high-quality renderings with fewer noise artifacts, as evidenced by the improved rendering quality of the floor in the fourth column and the chair in the sixth column. This demonstrates that better mapping performance leads to enhanced rendering quality. The joint optimization of semantic losses and photometric loss is mutually beneficial, facilitating more accurate global 3D reconstruction through semantic learning, which in turn improves the rendering performance of our Hier-SLAM++.
TABLE IV: Reconstruction performance Depth L1 (cm) on Replica. Best results are highlighted as first, second.
Methods Avg. R0 R1 R2 Of0 Of1 Of2 Of3 Of4
NICE-SLAM [33] 2.97 1.81 1.44 2.04 1.39 1.76 8.33 4.99 2.01
Vox-Fusion [43] 2.46 1.09 1.90 2.21 2.32 3.40 4.19 2.96 1.61
Co-SLAM [44] 1.51 1.05 0.85 2.37 1.24 1.48 1.86 1.66 1.54
ESLAM [35] 0.95 0.73 0.74 1.26 0.71 1.02 0.93 1.03 1.18
SNI-SLAM [11] 0.77 0.55 0.58 0.87 0.55 0.97 0.89 0.75 0.97
SGS-SLAM [23] 0.36
SemGauss-SLAM [24] 0.50 0.54 0.46 0.43 0.29 0.22 0.51 0.98 0.56
Hier-SLAM [25] 0.49 0.58 0.40 0.40 0.29 0.19 0.51 0.95 0.57
[…]
Hier-SLAM++ (binary) 0.49 0.59 0.40 0.40 0.30 0.17 0.52 0.95 0.57
IV-E Running time
The runtime results for the compared methods are provided in Tab. VI, covering both NeRF-based and Gaussian Splatting-based SLAM methods. Since the compared baseline using a flat representation (Hier-SLAM++-flat) cannot run successfully on an NVIDIA RTX 4090 due to GPU memory limitations, we report the runtime separately for RTX 4090 and L40S. In the first block of Tab. VI, we observe that, compared to NeRF-based methods [33], Gaussian-based state-of-the-art methods [16] achieve significantly higher operating speeds, benefiting from the fast rendering and efficient optimization capabilities inherent to the new 3D representation, i.e., Gaussian Splatting. Furthermore, with the carefully optimized forward and backward Gaussian rendering module introduced in our system, which serves as the core component of Gaussian Splatting-based SLAM systems, our method (Hier-SLAM++- w/o sem) achieves up to 2.4× faster tracking and 2.2× faster mapping compared to the best-performing method [16].
In the second block of Tab. VI, when integrating semantic information, our method maintains high efficiency by leveraging hierarchical semantic coding (Hier-SLAM++-one-hot), achieving nearly 3× faster tracking compared to flat semantic coding (Hier-SLAM++-flat). Compared to the state-of-the-art method [25], our Hier-SLAM++ attains similar operational speed but significantly outperforms in semantic segmentation accuracy. Thanks to the improved semantic decoder in our hierarchical semantic loss, our method achieves a $13\%$ mIoU improvement while incurring only an additional $0.04$ s per frame for tracking and $2$ s per frame for mapping. Further details on this improvement will be provided in the next section. Additionally, our binary version (Hier-SLAM++-binary) is able to achieve more efficient operation performance. Notably, our Hier-SLAM++ achieves a rendering speed of 2000 FPS, which increases to 3000 FPS when semantic information is not included.
TABLE V: Rendering performance PSNR, SSIM, LPIPS on Replica. Best results are highlighted as first, second.
Methods Metrics Avg. room0 room1 room2 office0 office1 office2 office3 office4
Visual SLAM
NICE-SLAM [33] PSNR $\uparrow$ 24.42 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94
SSIM $\uparrow$ 0.809 0.689 0.757 0.814 0.874 0.886 0.797 0.801 0.856
LPIPS $\downarrow$ 0.233 0.330 0.271 0.208 0.229 0.181 0.235 0.209 0.198
Vox-Fusion [43] PSNR $\uparrow$ 24.41 22.39 22.36 23.92 27.79 29.83 20.33 23.47 25.21
SSIM $\uparrow$ 0.801 0.683 0.751 0.798 0.857 0.876 0.794 0.803 0.847
LPIPS $\downarrow$ 0.236 0.303 0.269 0.234 0.241 0.184 0.243 0.213 0.199
Co-SLAM [44] PSNR $\uparrow$ 30.24 27.27 28.45 29.06 34.14 34.87 28.43 28.76 30.91
SSIM $\uparrow$ 0.939 0.910 0.909 0.932 0.961 0.969 0.938 0.941 0.955
LPIPS $\downarrow$ 0.252 0.324 0.294 0.266 0.209 0.196 0.258 0.229 0.236
ESLAM [35] PSNR $\uparrow$ 29.08 25.32 27.77 29.08 33.71 30.20 28.09 28.77 29.71
SSIM $\uparrow$ 0.929 0.875 0.902 0.932 0.960 0.923 0.943 0.948 0.945
LPIPS $\downarrow$ 0.239 0.313 0.298 0.248 0.184 0.228 0.241 0.196 0.204
SplaTAM [16] PSNR $\uparrow$ 34.11 32.86 33.89 35.25 38.26 39.17 31.97 29.70 31.81
SSIM $\uparrow$ 0.968 0.978 0.969 0.979 0.977 0.978 0.969 0.949 0.949
LPIPS $\downarrow$ 0.102 0.072 0.103 0.081 0.092 0.093 0.102 0.121 0.152
Semantic SLAM
SNI-SLAM [11] PSNR $\uparrow$ 29.43 25.91 28.17 29.15 31.85 30.34 29.13 28.75 30.97
SSIM $\uparrow$ 0.921 0.884 0.900 0.921 0.935 0.925 0.930 0.932 0.936
LPIPS $\downarrow$ 0.237 0.307 0.292 0.265 0.185 0.211 0.230 0.209 0.198
SGS-SLAM [23] PSNR $\uparrow$ 34.66 32.50 34.25 35.10 38.54 39.20 32.90 32.05 32.75
SSIM $\uparrow$ 0.973 0.976 0.978 0.981 0.984 0.980 0.967 0.966 0.949
LPIPS $\downarrow$ 0.096 0.070 0.094 0.070 0.086 0.087 0.101 0.115 0.148
SemGauss-SLAM [24] PSNR $\uparrow$ 35.03 32.55 33.32 35.15 38.39 39.07 32.11 31.60 35.00
SSIM $\uparrow$ 0.982 0.979 0.970 0.987 0.989 0.972 0.978 0.972 0.978
LPIPS $\downarrow$ 0.062 0.055 0.054 0.045 0.048 0.046 0.069 0.078 0.093
Hier-SLAM [25] PSNR $\uparrow$ 35.70 32.83 34.68 36.33 39.75 40.93 33.29 32.48 35.33
SSIM $\uparrow$ 0.980 0.976 0.979 0.987 0.988 0.989 0.975 0.971 0.976
LPIPS $\downarrow$ 0.067 0.060 0.063 0.052 0.050 0.049 0.083 0.081 0.094
Hier-SLAM++ (one-hot) PSNR $\uparrow$ 35.61 32.76 34.64 36.27 39.65 40.12 33.46 32.49 35.52
SSIM $\uparrow$ 0.980 0.978 0.979 0.987 0.987 0.985 0.977 0.971 0.977
LPIPS $\downarrow$ 0.067 0.058 0.058 0.051 0.053 0.064 0.077 0.082 0.100
[…]
SSIM $\uparrow$ 0.981 0.975 0.979 0.988 0.987 0.990 0.978 0.971 0.976
LPIPS $\downarrow$ 0.066 0.064 0.061 0.049 0.056 0.043 0.076 0.081 0.098
TABLE VI: Runtime on Replica/R0. Best results are highlighted as first.
Methods Tracking Mapping Tracking Mapping
/Iteration /Iteration /Frame /Frame
RTX 4090 NICE-SLAM [33] 122.42 104.25 1.22 6.26
SplaTAM [16] 44.27 50.07 1.77 3.00
Hier-SLAM++ (w/o sem) 18.71 22.93 0.75 1.38
Hier-SLAM++ (one-hot) 37.63 194.78 1.50 11.69
L40S Hier-SLAM [25] 61.23 170.30 2.45 10.22
Hier-SLAM++ (flat) 168.94 204.25 6.75 12.26
Hier-SLAM++ (one-hot) 62.21 212.62 2.49 12.76
Hier-SLAM++ (binary) 58.91 210.65 2.36 12.64
  • Hier-SLAM++ (w/o sem) represents our proposed system without semantic information.
  • The units are as follows: Tracking/Iteration (ms), Mapping/Iteration (ms), Tracking/Frame (s), and Mapping/Frame (s).
TABLE VII: Semantic performance mIoU (%) and Parameter usage (MB) on Replica. Evaluations are conducted over a total of 102 semantic classes. Results are highlighted as first, second.
Methods Avg. R0 R1 R2 Of0
mIoU (%) NIDS-SLAM [12] 82.37 82.45 84.08 76.99 85.94
DNS-SLAM [7] 84.77 88.32 84.90 81.20 84.66
Hier-SLAM [25] 76.44 76.62 78.31 80.39 70.43
Hier-SLAM++ (flat) 90.35 91.21 90.62 89.11 90.45
Hier-SLAM++ (one-hot) 89.41 86.38 89.26 91.55 90.46
Hier-SLAM++ (binary) 81.27 82.77 73.87 89.64 78.80
SGS-SLAM [23] 92.72 92.95 92.91 92.10 92.90
SemGauss-SLAM [24] 96.34 96.30 95.82 96.51 96.72
SNI-SLAM [11] 87.41 88.42 87.43 86.16 87.63
Param (MB) Hier-SLAM [25] 910.50 793 1126 843 880
Hier-SLAM++ (flat) 2662.25 2355 3072 2560 2662
Hier-SLAM++ (one-hot) 927.50 814 1126 867 903
Hier-SLAM++ (binary) 591.75 528 690 563 586
  • The gray background indicates that the different evaluation methods are used, where evaluations are conducted on a subset of the total classes.
IV-F Hierarchical semantic understanding
We perform semantic understanding experiments on the synthetic dataset Replica [41] to demonstrate the comprehensive performance of our proposed method. Replica [41] is a synthetic indoor dataset that includes 102 semantic classes with high-quality semantic annotations. Similar to previous works [24, 7, 11], we evaluate the mIoU performance as the quantitative semantic results in Tab. VII, where mIoU (%) is reported by comparing the rendered semantic map with the ground truth across all original semantic classes (102 classes). To further analyze the storage requirements of semantic SLAM methods, we report the parameter storage usage in the second block of Tab. VII.
From Tab. VII, our method, along with the flat coding version and the proposed one-hot version, both demonstrate superior semantic performance compared to all existing methods. By utilizing hierarchical semantic representation, Hier-SLAM++ (one-hot) achieves a similar semantic performance with less than a $1\%$ difference in mIoU while requiring 4.5× less storage compared to the flat version. The training time shown in Tab. VI demonstrates that our hierarchical version (Hier-SLAM++-one-hot) requires only 37% of the tracking time per frame while maintaining similar frame mapping performance compared to the flat version (Hier-SLAM++-flat). The comparable mIoU performance, along with significantly reduced storage and computational cost, clearly demonstrates the effectiveness and efficiency of our proposed method. Furthermore, our binary version (Hier-SLAM++-binary) achieves state-of-the-art semantic understanding while achieves the best parameter storage efficiency, reducing storage usage to 63% of the one-hot version and requiring only 20% of the storage compared to the flat version. This reduction in storage usage highlights the efficiency of our proposed method. Because SGS-SLAM [23], SemGauss-SLAM [24] and SNI-SLAM [11] use different evaluation methods that restrict the mIoU calculation to a subset of the global semantic classes. Specifically, SGS-SLAM computes the average IoU for the semantic classes visible in each frame, and SemGauss-SLAM together with SNI-SLAM [11] evaluates performance based on 52 visible classes across all evaluation sequences, making direct comparisons unfair.
[…]
Overall, our method demonstrates superior performance in semantic understanding while significantly reducing storage requirements and training time, benefiting from the proposed hierarchical semantic representation.
Figure 4: Visualization of our semantic rendering performance on the Replica [41] dataset. The first four rows demonstrate rendered semantic segmentation in a coarse-to-fine manner. The fifth row exhibits the finest semantic rendering, equivalent to the flat representation with $102$ original semantic classes from the Replica dataset. The last row visualizes the semantic ground truth for comparison.
IV-G Scaling up capability
To showcase the scaling-up capability of Hier-SLAM++, we evaluate it on the real-world complex dataset, ScanNet [40], which contains up to 550 distinct semantic classes. In contrast to Replica [41], where the semantic ground truth is synthesized from a global world model and can be considered ideal, the semantic annotations in ScanNet are significantly noisier, containing considerable noise and incorrect or unclear class boundaries. Moreover, the dataset contains noisy depth sensor inputs and blurred color images, making semantic understanding particularly challenging in this scenes.
Using the flat semantic representation cannot even run successfully due to memory limitations. In contrast, we establish the hierarchical tree with the assistance of LLMs and 3D Generative Models, as introduced in Section-III, which guides the compaction of the coding from the original 550 semantic classes to 73 semantic symbolic codings, resulting in over 7 times reduction in coding usage. As visualized in Fig. 5, our estimated 3D global semantic map at different levels demonstrates a coarse-to-fine semantic understanding, showcasing our method’s scaling-up capability in handling real-world complex scenes.
TABLE VIII: Monocular quantitative results on the TUM RGB-D dataset. Best results are highlighted as first, second.
Sequences fr1-desk fr2-xyz
Loop Method RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/ DROID-SLAM [48] 1.80 - - - 0.50 - - -
ORB-SLAM2 [29] 2.00 - - - 0.60 - - -
Photo-SLAM [19] 1.54 20.97 0.743 0.228 0.98 21.07 0.726 0.166
[…]
MonoGS [17] 4.15 21.02 0.703 0.362 4.79 22.21 0.728 0.288
Hier-SLAM++ 6.20 21.91 0.836 0.269 3.22 24.67 0.946 0.110
IV-H Monocular performance
We present our monocular SLAM performance in Tab. VIII. For camera tracking, our method achieves competitive performance compared to state-of-the-art approaches without loop closure. Specifically, compared to the state-of-the-art Gaussian Splatting-based methods [17], our method exhibits similar tracking performance while achieving superior rendering quality. For SLAM methods with loop closure, which typically demonstrate superior camera tracking performance due to well-executed loop closure mechanisms that significantly enhance tracking accuracy and overall SLAM performance, our method still achieves superior rendering accuracy compared to Photo-SLAM [19], which incorporates loop closure. This highlights the strong rendering capabilities of our proposed method. With the assistance of the geometric prior provided by the feed-forward method, the joint optimization of both photometric loss and aligned depth loss contributes to improved rendering performance in the whole SLAM system.
Figure 5: Visualization of the established semantic 3D map across multiple levels, illustrating a coarse-to-fine semantic understanding of the complex scene. The bottom of the figure presents localization, rendering, and depth performance, offering a comprehensive overview of Hier-SLAM++’s effectiveness and demonstrating the scaling-up capability of our proposed method.
IV-I Ablation study
We evaluate SLAM performance under different tree designs in Tab. IX. The evaluation is conducted using the baseline method, Hier-SLAM [25], ensuring that any performance differences are solely attributable to the tree structures. In Tab. IX, each metric represents the average result from two runs on the first four sequences—Room0, Room1, Room2, and Office0—from the Replica dataset. We generate two different 5-level trees using our proposed tree generation procedure for the Replica dataset and incorporate them into the semantic SLAM system. As shown in Tab. IX, all methods demonstrate good performance, which can be attributed to the well-designed tree structures generated through both LLMs and 3D Generative Models. Additionally, the fully optimized hierarchical semantic loss contributes to the robustness of SLAM performance. Among all compared methods, tree-1 achieves the best semantic mIoU performance, the highest depth mapping accuracy, and close rendering quality (PSNR, SSIM, and LPIPS) compared with other settings. Therefore, we select tree-1 as the hierarchical tree structure leveraged within our semantic SLAM system.
TABLE IX: Performance of different trees on Replica. Best results are highlighted in bold.
[…]
Hier-SLAM [25] 75.517 0.292 0.416 35.885 0.982 0.057
[…]
Hier-SLAM++${}^{\text{tree2}}$ 75.061 0.280 0.413 35.792 0.982 0.057
V Conclusion
We introduced Hier-SLAM++, a neuro-symbolic semantic 3D Gaussian Splatting SLAM system that supports both RGB-D and monocular inputs, incorporating a novel hierarchical categorical representation. Specifically, we proposed a compact and generalized hierarchical representation that encodes both semantic meaning and geometric attributes (i.e., size and shape), leveraging the capabilities of LLMs and 3D generative models. To enhance global semantic understanding, we introduced an improved hierarchical semantic loss. Furthermore, we extended the system from an RGB-D-only SLAM to support monocular input, utilizing geometric priors derived from a feed-forward SLAM approach. Experimental results demonstrate that Hier-SLAM++ achieves superior or on-par performance with state-of-the-art NeRF-based and Gaussian-based SLAM methods in terms of tracking, mapping, and semantic understanding accuracy, while significantly reducing storage requirements and runtime. These advances establish Hier-SLAM++ as a robust solution for semantic 3D understanding across diverse environments.
Figure 6: Visualization of RGB image rendering performances on the Replica dataset [41]. The first row presents the image rendering results of our Hier-SLAM++. The second to fourth rows show the rendering outputs of the comparison methods. The last row displays the ground truth images for reference. We use red boxes to highlight key differences regions within the image. The visual comparison clearly demonstrates that our method achieves superior rendering quality, closely resembling the ground truth images compared to state-of-the-art approaches.
APPENDIX
I Appendix 1. Implementation Details for Tree Generation
During the hierarchical tree generation process, we utilize LLMs to group flat semantic classes at different levels based on physical size, semantic function, and to summarize and refine geometric grouping, ensuring balanced, meaningful, and descriptive outputs. After generating each tree level, we further employ LLMs as validators to examine the results at each level, ensuring that all input categories are fully included with no omissions and that no extra categories are mistakenly introduced.
Size Grouping: The prompt used for the initial grouping based on object physical sizes is illustrated as follows:
Code I.1: Prompt for Size Grouping You are a smart assistant tasked with dividing the following items into meaningful groups based on their properties. The key requirements are:
Balance: The number of items in each group should be as evenly distributed as possible. A difference of 1 item between groups is acceptable, but larger differences should be avoided.
Meaningfulness: The groups must be meaningful, based on the items’ inherent characteristics. Grouping should make logical sense to a human observer.
Descriptive Group Names: Each group must have a clear and descriptive name that reflects its characteristics.
Goal: Cluster items into size groups, based on their typical physical size in scenes. The groups must strictly be by sizes, for example: small, small medium, medium, large, extra large, etc.
Items: {classes input}
Ensure groups are meaningful and provide descriptive group names. Output must follow this JSON format:
    "<GROUP_1>": ["<ITEM_1>", ...],
    "<GROUP_2>": ["<ITEM_2>", ...],
Function Grouping: The prompt adopted for the subsequent semantic function grouping is as follows:
Code I.2: Prompt for Function Grouping You are a smart assistant tasked with dividing the following items into meaningful groups based on their properties. The key requirements are:
Balance: The number of items in each group should be as evenly distributed as possible. A difference of 1 item between groups is acceptable, but larger differences should be avoided.
Meaningfulness: The groups must be meaningful, based on the items’ inherent characteristics. Grouping should make logical sense to a human observer.
Descriptive Group Names: Each group must have a clear and descriptive name that reflects its characteristics.
Goal: Cluster items in the ’size’ size group into functionality-based groups. The name of the clusters should not be too specific, it could be as general like storage, furniture, etc. Ensure that items in each group serve similar purposes or have similar functionalities.
Items: {classes input}
Ensure groups are meaningful and provide descriptive group names. Output must follow this JSON format:
    "<GROUP_1>": ["<ITEM_1>", ...],
    "<GROUP_2>": ["<ITEM_2>", ...],
Summarize and Refine for Geometric Grouping: After grouping with the 3D Generative Model, we employ LLMs to summarize the clustering results with descriptive labels and balance the nodes within each group. The prompt for LLMs is illustrated as follows:
Code I.3: Prompt for Geometric Grouping Results The following groups of indoor scene items have been clustered based on their shape: {formatted_clusters}
Instructions: Balance: Ensure the groups are as evenly distributed as possible. A difference of 1 item between groups is acceptable. If needed, move a small number of items from one group to another to achieve balance, feel free to even remove a group if it has only one member (just dont leave any groups empty), ensuring that the groups remain meaningful.
Meaningful Naming: After balancing the groups, assign a descriptive and meaningful name to each group, based on the shared shape characteristic. The name should clearly reflect the shape or geometric property of the items in that group. Always prefer singular shapes names like box, rectangle, soft, flat, etc.
No Dublications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
The group names must strictly by shapes and DO NOT leave any group empty, you could remove it if its empty.
The output must be the same JSON format as below:
    "<GROUP_1>": ["<ITEM_1>", ...],
    "<GROUP_2>": ["<ITEM_2>", ...],
LLM Validators: The prompt used for LLM validator is as following:
Code I.4: Prompt for LLM validator The following groups of items in a scene have been clustered based on their shape: {formatted_clusters}
Instructions: Balance: Ensure the groups are as evenly distributed as possible. A difference of 1 item between groups is acceptable. If needed, move a small number of items from one group to another to achieve balance, remove a group if it has only one member, ensuring that the groups remain meaningful.
Meaningful Naming: After balancing the groups, assign a descriptive and meaningful name to each group, based on the shared shape characteristic. The name should clearly reflect the shape or geometric property of the items in that group. Always prefer singular shapes names like box, rectangle, soft, flat, etc.
New Group: If it is necessary to create a new group, feel free to do so, DO NOT create any non-original items.
No Dublications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
The output must be the same JSON format as below:
    "<GROUP_1>": ["<ITEM_1>", ...],
    "<GROUP_2>": ["<ITEM_2>", ...],

### v2

@@
 The parameter usage in semantic SLAM systems increases significantly with the growing complexity of the environment, making scene understanding particularly challenging and costly.
-To address this problem, we introduce a novel and general hierarchical representation that encodes both semantic and geometric information in a compact form into 3D Gaussian Splatting, leveraging the capabilities of large language models (LLMs) as well as the 3D generative model.
+To address this problem, we introduce a novel hierarchical representation that encodes both semantic and geometric information in a compact form into 3D Gaussian Splatting, leveraging the capabilities of large language models (LLMs) as well as the 3D generative model.
 By utilizing the proposed hierarchical tree structure, semantic information is symbolically represented and learned in an end-to-end manner.
-We further introduce a novel semantic loss designed to optimize hierarchical semantic information through both inter-level and cross-level optimization.
+We further introduce an advanced semantic loss designed to optimize hierarchical semantic information through both Intra-level and Inter-level optimizations.
 Additionally, we propose an improved SLAM system to support both RGB-D and monocular inputs using a feed-forward model.
 To the best of our knowledge, this is the first semantic monocular Gaussian Splatting SLAM system, significantly reducing sensor requirements for 3D semantic understanding and broadening the applicability of semantic Gaussian SLAM system.
-We conduct experiments on both synthetic and real-world datasets, demonstrating superior or on-par performance with state-of-the-art NeRF-based and Gaussian-based SLAM systems, while significantly reducing storage and training time requirements.
+We conduct experiments on both synthetic and real-world datasets, demonstrating superior or on-par performance with state-of-the-art methods, while significantly reducing storage and training time requirements.
+Our project page is available at: https://hierslampp.github.io/.
 Index Terms:
@@
 I Introduction
-Visual Simultaneous Localization and Mapping (SLAM) is a key technology for ego-motion estimation and scene perception, widely employed in applications across various domains, such as autonomous drones [1], self-driving vehicles [2], and interactive experiences in Augmented Reality (AR) and Virtual Reality (VR) [3].
-Semantic information, which provides high-level knowledge about an environment, is indispensable for comprehensive scene understanding.
-Thus, it plays a crucial role in enabling intelligent robots to perform various complex tasks.
-In recent decades, advances in image segmentation and enhanced map representations have accelerated progress in semantic visual SLAM [4, 5, 6, 7, 8, 9, 10, 11, 7, 12].
+Visual Simultaneous Localization and Mapping (SLAM) is a key technology for ego-motion estimation and scene perception, and is widely employed across various domains such as autonomous drones [1], self-driving vehicles [2], and interactive experiences in Augmented Reality (AR) and Virtual Reality (VR) [3].
+Semantic information, which provides high-level knowledge about the environment, is indispensable for comprehensive scene understanding.
+It plays a crucial role in enabling intelligent robots to perform complex tasks.
+In recent decades, advances in image segmentation and enriched map representations have significantly accelerated progress in semantic visual SLAM [4, 5, 6, 7, 8, 9, 10, 11, 12].
 Figure 1: (a).
@@
 The rendered semantic map at different levels shows a coarse-to-fine understanding, beneficial for real-world scenarios with shifting perspectives from distant to close.
-Recently, 3D Gaussian Splatting has gained attention as a leading 3D world representation method [13, 14, 15], thanks to its fast rendering and optimization efficiency enabled by the highly parallelized rasterization of 3D primitives.
-Specifically, 3D Gaussian Splatting models the continuous distributions of geometric parameters through Gaussian distributions.
-This approach not only improves the quality of 3D representation but also supports efficient optimization, making it particularly well-suited for SLAM tasks.
+Recently, 3D Gaussian Splatting has emerged as a powerful representation for 3D scenes [13, 14, 15], owing to its fast rendering and optimization efficiency enabled by the highly parallelized rasterization of 3D primitives.
+Specifically, it models the continuous distribution of geometric attributes using Gaussian functions.
+This formulation not only enhances the 3D representation fidelity but also facilitates efficient optimization, making it particularly suitable for SLAM tasks.
 While promising, current SLAM systems based on 3D Gaussian Splatting [16, 17, 18, 19, 20] primarily focus on geometric reconstruction.
-However, the lack of semantic integration limits its potential for complex tasks such as visual navigation, planning, and autonomous driving.
-Addressing this gap by incorporating semantic understanding without sacrificing efficiency remains a key challenge.
-A naive solution would be to augment 3D primitives with discrete semantic labels and parameterize their distribution using a categorical distribution, i.e., a flat Softmax-based representation.
-However, 3D Gaussian Splatting is already a storage-intensive approach [21, 22], requiring a large number of 3D primitives with multiple parameters to achieve realistic rendering.
-Adding semantic distribution parameters would further increase storage requirements and processing time, with grow linear in the number of semantic classes, making it particularly impractical for complex scene understanding.
-Furthermore, given a fixed number of observations (video frames), increasing the number of optimizable parameters leads to the curse of dimensionality, resulting in degraded optimization performance in high-dimensional spaces, such as when handling a large number of semantic classes.
-Overall, this added overhead for semantic learning is especially problematic for robotics applications, where real-time performance and limited resources are critical.
-Recent studies have explored alternative methods to mitigate these challenges.
-For instance, [23] directly learns a 3-channel RGB visualization to avoid semantic labels learning.
-[24] adopts a flat semantic representation and relies on a pre-trained foundation model to generate an embedding feature map rather than an explicit 3D semantic label map.
-Unlike flat representations, semantic information naturally organizes into a hierarchical structure of classes, as shown in Fig.
-1.
-The attributes (such as structures) and properties (such as surfaces) of the semantic classes are extracted as symbolic nodes to generate a hierarchical tree, where each semantic class is represented as a hierarchical expression based on a root-to-leaf path.
-This hierarchy enable efficient coding of extensive information with a relatively small number of symbolic nodes to learn, resulting in a compact representation.
-For example, a binary tree with a depth of $10$ can represent $2^{10}$ classes, allowing for the encoding of 1,024 classes using only 20 codes (i.e., $2\times 10$ through 2-dimensional Softmax coding at each level), enabling a reduction of up to $O(\log N)$.
-Building upon this concept, Hier-SLAM [25] was first introduced as an RGB-D Gaussian Splatting SLAM that employs a hierarchical tree to represent semantic information in an unknown environment, achieving strong performance with improved storage efficiency and operational speed.
-Hier-SLAM [25] have demonstrated the advantages and potential of incorporating hierarchical categorization for a semantic RGB-D SLAM system.
-However, it is limited by a simplistic hierarchical tree representation and lacks robust semantic optimization.
-To address these limitations and further improve its effectiveness, we propose Hier-SLAM++, a comprehensive neuro-symbolic semantic Gaussian Splatting SLAM framework that incorporates an advanced hierarchical representation for semantic information.
-Different from Hier-SLAM, we incorporate both geometric (i.e., shape and size details) and semantic information into the hierarchical symbolic tree expression using both Large Language Models (LLMs) and 3D Generative models.
-Specifically, we utilize a text-to-3D generative model [26] to extract shape information for each class and organize it into hierarchical levels.
-Additionally, we leverage LLMs to include both semantic and size information in the tree structure, effectively addressing the lack of size information in 3D generative models.
-With the established tree, semantic information in the environment is represented as symbolic nodes along root-to-leaf paths, to be learned in an end-to-end manner during the SLAM operation.
-To enhance semantic understanding, we introduce an improved hierarchical loss function that incorporates both inter-level and cross-level optimizations.
-This coarse-to-fine hierarchical approach based on the established hierarchical representation is beneficial for real-world applications, especially those involving observations from distant to nearby views.
-Furthermore, our SLAM system is extended to support both RGB-D and monocular inputs.
-We utilize the geometric output of the 3D feed-forward model [27] as a geometric prior, removing the depth dependency of existing SLAM methods.
-This enables Hier-SLAM++ to perform semantic SLAM without relying on dedicated depth sensors, broadening the range of applicable scenarios.
-Finally, we enhance and refine the 3D Gaussian Splatting SLAM to improve both performance and efficiency.
-We conduct experiments on multiple synthetic and real-world datasets.
-The results demonstrate that our SLAM system outperforms or is on par with state-of-the-art NeRF-based and Gaussian-based SLAM methods.
-Additionally, our approach effectively reduces storage and training time.
-The technical contributions of our work include:
-1) We introduce an advanced hierarchical tree representation that integrates semantic, shape, and size information, leveraging the capabilities of both LLMs and 3D generative models.
-This tree-based coding efficiently compresses semantic information while maintaining its inherent hierarchical structure, resulting in reduced memory usage and shorter training times.
-2) We present a novel optimization approach for the proposed semantic hierarchical representation, integrating both inter-level and cross-level losses to ensure thorough optimization across all levels of the hierarchical semantic information.
-3) We propose an improved semantic SLAM system capable of supporting both RGB-D and monocular inputs by leveraging a 3D feed-forward model.
-To the best of our knowledge, Hier-SLAM++ is the first Gaussian Splatting based semantic SLAM system to support monocular input, significantly reducing sensor requirements and broadening its application scope.
-4) We conduct extensive experiments on both synthetic and real-world datasets, comprehensively demonstrating the effectiveness of our method and its improvements.
+However, the absence of semantic integration limits their applicability in complex downstream tasks that require high-level scene understanding.
+To enable semantic understanding in SLAM systems, a straightforward approach, analogous to the geometric modeling of 3D Gaussians, is to augment each 3D primitive with a discrete semantic label and represent its distribution using a categorical distribution, typically implemented via a flat Softmax-based representation.
+However, 3D Gaussian Splatting is already a storage-intensive approach [21, 22], requiring numerous primitives with rich parameters to achieve high-quality rendering.
+Adding semantic distributions for each primitive further increases memory and computational costs, with complexity growing linearly with the number of semantic classes, making it impractical for complex semantic understanding.
+Moreover, given a limited number of observations (e.g., video frames), introducing high-dimensional semantic parameters leads to the curse of dimensionality, hindering optimization performance.
+This added overhead is especially problematic in robotics, where computational efficiency and resource constraints are critical.
+Therefore, a key challenge is to design semantic encoding strategies that are both efficient and scalable, enabling robust semantic SLAM without compromising runtime and memory efficiency.
+Semantic information often exhibits a natural hierarchical structure.
+As illustrated in Fig.
+1, semantic classes can be abstracted into multi-level concepts, such as structures and surfaces, which can be further organized into a hierarchical symbolic tree.
+With careful design, this structure enables rich semantics to be encoded efficiently using a small number of symbolic nodes, yielding a compact and scalable representation.
+For instance, a binary tree of depth $10$ can represent up to $2^{10}$ classes using only $20$ nodes (i.e., $2\times 10$ nodes through 2-dimensional Softmax at each level), reducing the complexity from $O(N)$ to $O(\log N)$.
+Building upon this concept, we propose Hier-SLAM++, a comprehensive neural-symbolic semantic SLAM framework based on 3D Gaussian Splatting, featuring a novel hierarchical representation for semantic information.
+To design an effective hierarchical structure, we incorporate both semantic and geometric cues, including shape and size information.
+Specifically, we leverage a text-to-3D generative model [23] to extract shape-aware features for each class and organize them into hierarchical levels.
+Additionally, we employ Large Language Models (LLMs) to enrich the hierarchy with semantic relations and size priors, compensating for the limitations of generative models.
+With the constructed tree, semantic information in the environment is represented as symbolic nodes along a hierarchical path and learned end-to-end during SLAM optimization.
+Specifically, we propose two types of hierarchical semantic representations, one-hot and binary semantic embedding, which offer different degrees of compactness.
+To enhance semantic understanding, we introduce a novel hierarchical loss function that performs joint optimization at multiple levels, including Intra-level and Inter-level supervision.
+Furthermore, Hier-SLAM++ supports both RGB-D and monocular settings.
+We employ the geometric output of a 3D feed-forward model [24] as a depth prior, effectively removing the reliance on dedicated depth sensors.
+This enables monocular semantic SLAM and broadens the applicability of our system.
+We further refine the 3D Gaussian Splatting-based SLAM pipeline to improve both tracking performance and runtime efficiency.
+Experiments on both synthetic and real-world datasets demonstrate that Hier-SLAM++ achieves superior or on-par performance with state-of-the-art methods in localization, mapping, and semantic understanding under the RGB-D setting, while maintaining competitive performance in the monocular setting.
+The main contributions are as follows:
+1) We propose a hierarchical tree structure that integrates semantic, shape, and size information, constructed using LLMs and 3D generative models.
+This representation efficiently encodes semantic messages while preserving its structural hierarchy, leading to reduced memory consumption and faster training.
+2) Based on the generated hierarchical tree, we propose two types of semantic embeddings that offer different degrees of compactness.
+Furthermore, we introduce a novel optimization strategy for hierarchical semantic representations, incorporating both Intra-level and Inter-level losses to enable comprehensive refinement of both embeddings.
+3) We develop an enhanced SLAM framework that supports both RGB-D and monocular settings by leveraging a 3D feed-forward model as a geometric prior.
+To the best of our knowledge, Hier-SLAM++ is the first Gaussian Splatting-based semantic SLAM system capable of operating with monocular input, significantly lowering sensor requirements and improving deployment flexibility.
+This paper builds upon our previous work [25], with the following major extensions: 1) The tree construction process is enhanced by integrating 3D generative models and improving the use of LLMs, resulting in a more effective symbolic tree structure.
+2) In addition to the one-hot semantic representation from our earlier work, we propose a binary semantic representation for further compactness.
+3) We improve semantic optimization with a new hierarchical loss formulation and introduce a comprehensive hierarchical optimization strategy applicable to both one-hot and binary representations.
+4) We propose a monocular pipeline to enable robust monocular SLAM by leveraging 3D feed-forward models, thereby extending the RGB-D semantic SLAM system into a unified framework that supports both RGB-D and monocular inputs.
+Additionally, we include new real-world datasets and experiments to further validate the proposed approach.
 II Related Work
 II-A Visual SLAM system
-Visual Simultaneous Localization and Mapping (SLAM) using visual sensors is a long-standing and active research topic that focuses on the simultaneous estimation of ego-motion and the reconstruction of a global 3D map.
-Traditional SLAM methods [28, 29, 30, 31], developed over many years, have demonstrated both high accuracy and robustness, even in challenging environments.
-This makes them an indispensable frontend foundation for various real-world applications, such as autonomous driving and drones.
-In recent years, several new 3D representations have emerged in the 3D computer vision fields.
-With the advent of neural implicit representations, numerous works [32, 33, 34, 35] have demonstrated the effectiveness of NeRF-based SLAM approaches.
-These new representations introduce novel capabilities to the SLAM community, including enhanced rendering capability.
-3D Gaussian Splatting is the latest advancement in the 3D vision field, offering significantly faster rendering speed and strong performance compared with neural implicit representations.
-Recent works [16, 17, 20, 19] have demonstrated new SLAM capabilities by integrating 3D Gaussian Splatting techniques, further pushing the boundaries of real-time scene understanding and reconstruction.
+Visual Simultaneous Localization and Mapping (SLAM) is a long-standing topic in computer vision and robotics.
+It jointly estimates camera poses and reconstructs global maps by optimizing appearance-based objectives [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
+One classical monocular-based approach aims to estimate accurate camera poses and a sparse global map, by minimizing the reprojection error terms [28, 29, 30], or by minimizing the photometric losses [31, 32, 33].
+DTAM [34] takes a different direction by performing dense reconstruction and camera tracking with RGB inputs, though real-time capability relys on GPU support.
+Another category of methods adopts RGB-D input for dense 3D reconstruction, which is able to provide richer information.
+The works [35, 36] jointly estimate camera poses and dense surfaces from RGB-D input using a volumetric TSDF map.
+Alongside advances in purely geometric SLAM, semantic SLAM has gained increasing attention for its ability to provide high-level scene understanding [37, 38, 39, 40, 41, 42, 43, 44, 8, 6].
+Sparse semantic SLAM methods rely on pre-scanned models [37, 38] or use bounding boxes [39, 40, 41] and quadrics [42], though accuracy remains limited.
+In contrast, semantic understanding in dense SLAM becomes more tractable and effective when combined with dense observations from RGB-D inputs [45, 46, 47] or dense visual SLAM pipelines [44].
+More recently, new paradigms have emerged to enhance dense SLAM with richer scene understanding across both monocular and RGB-D inputs, such as implicit neural fields and 3D Gaussian Splatting.
+Furthermore, semantic integration allows current SLAM systems to move beyond geometric reconstruction to deliver high-level scene interpretation, such as class-level mapping, semantic navigation, and task-aware perception.
+This significantly broadens its applications in real-world fields.
+Some works, such as Kimera [48], adopt dynamic scene graphs for representing entities within the observed environment  [49, 50, 51, 52, 53].
+In contrast, we propose a generalizable hierarchical symbolic representation for modeling semantics in the world, leveraging large language models and 3D generative priors, along with symbolic capabilities [54, 55] to abstract semantic concepts without relying on scene-specific definitions.
+Crucially, our symbolic hierarchy is carefully designed to produce compact semantic embeddings and is seamlessly integrated into the neural learning process within the SLAM system, offering both structural abstraction and interpretability for semantic encoding.
 II-B Neural implicit visual SLAM
-Neural Radiance Fields (NeRF) [36] is a neural rendering technique that represents 3D scenes using a fully connected neural network.
-iMap [32] proposes a real-time keyframe-based RGB-D SLAM method that uses a multilayer perceptron (MLP) as the sole scene representation.
-Nice-SLAM [33] introduces a dense SLAM system that incorporates multi-level local information by employing a hierarchical scene representation.
-Nicer-SLAM [34] extends this approach by proposing a monocular SLAM system with a hierarchical neural implicit map representation, eliminating the dependency on depth information.
-Integrating SLAM with semantic understanding, many works [7, 11, 12] have explored neural implicit representations for semantic mapping and localization tasks.
-DNS-SLAM [7] incorporates 2D semantic priors with a coarse-to-fine geometry representation to embed semantic information into the map.
-SNI-SLAM [11] integrates geometry, appearance, and semantic features into a shared feature space to enhance the robustness of the SLAM method.
-However, these approaches are constrained by the slow convergence of neural implicit map representations, which often leads to performance degradation when coupled with semantic objectives [37, 38].
-In contrast, Gaussian Splatting provides significant advantages, including fast rendering performance and high-density reconstruction quality, making it a promising alternative for map representation.
+Neural Radiance Fields (NeRF) [56] introduce a novel scene representation paradigm, inspiring the integration of NeRF with SLAM pipelines for simultaneous localization and high-quality mapping.
+Unlike traditional SLAM methods that rely on explicit geometry, NeRF-SLAM models represent scenes implicitly, enabling continuous and photorealistic reconstruction from sequential views.
+Early work such as iNeRF [57] explored solely pose tracking with a fixed pre-trained neural field.
+iMap [58] advanced this idea by proposing a keyframe-based RGB-D SLAM that jointly optimizes camera poses and a global implicit map using a single MLP.
+NICE-SLAM [59] enhances robustness and scalability by introducing a hierarchical neural implicit representation, which is later extended to the monocular setting in NICER-SLAM [60].
+Recent works focus on improving efficiency: ESLAM [61] leverages a hybrid representation combining implicit TSDFs with multi-scale feature planes to achieve faster operation; Point-SLAM [62] adopts a point-based neural representation to enable efficient and detail-preserving dense SLAM.
+To further enhance scene understanding beyond geometry and appearance, researchers also incorporate semantics into NeRF-based SLAM systems [7, 11, 12, 63].
+Some works leverage the semantic information extracted from 2D observations to integrate them into the visual SLAM system and generate the global semantic map [7, 63].
+SNI-SLAM [11] proposes a semantic RGB-D SLAM which learn semantics by integrating geometry, appearance, and semantic features into a shared feature space.
+However, these methods remain constrained by the slow convergence and high overhead of neural implicit representations [64, 65], especially when jointly optimizing semantics.
+The inherent inefficiency of NeRF representations still conflicts with the real-time demands of SLAM systems—let alone the added cost from semantic integration.
+In contrast, Gaussian Splatting provides significant advantages, including fast rendering performance and high-density reconstruction quality, making it a promising new kind of map representation.
 II-C 3D Gaussian Splatting SLAM
-3D Gaussian Splatting has recently gained attention as a promising 3D representation due to its fast rendering performance and high-density reconstruction quality.
-Leveraging this 3D representation, SplaTAM [16] uses silhouette guidance for pose estimation and map reconstruction in RGB-D SLAM systems, while MonoGS [17] supports both monocular and RGB-D SLAM.
-Photo-SLAM [19] utilizes 3D points generated by ORB-SLAM as position parameters for Gaussian Splatting primitives within the SLAM system.
-These studies highlight the strong potential of 3D Gaussian Splatting in SLAM tasks across various scenarios [16, 17, 18, 19, 20].
-However, integrating semantic understanding into SLAM presents a significant challenge, as it requires jointly solving three high-dimensional optimization problems with varying value ranges and convergence behaviors, while also significantly increasing storage requirements.
-Focus on the integration of semantic information into the Gaussian Splatting SLAM, SGS-SLAM [23] added RGB 3-channels to create a semantic visualization map to avoid true semantic labels learning.
-SemGauss-SLAM [24] uses a flat semantic representation supervised by a large pre-trained foundation model.
-However, these methods overlook the natural hierarchical structure of the real world.
-Additionally, relying on large foundation models increases the complexity of the neural network and its computational demands, with performance largely dependent on the embeddings from these pre-trained models.
-Eliminating the dependency on foundation models, this work proposes a simple yet effective hierarchical representation for semantic understanding in the SLAM system which largely reduces the training storage and time.
+3D Gaussian Splatting has recently emerged as a promising 3D representation due to its fast rendering and high-fidelity reconstruction.
+Compared to implicit representations, it models the environment as a collection of explicit Gaussian distributions with geometric and appearance attributes.
+Several SLAM systems have adopted this representation in their pipelines.
+SplaTAM [16] proposes a RGB-D SLAM framework with isotropic Gaussians representation using silhouette guidance for pose estimation, while MonoGS [17] presents an efficient solution supporting both monocular and RGB-D inputs.
+Photo-SLAM [19] uses 3D points from ORB-SLAM to initialize Gaussian positions; GS-SLAM [18] introduces a coarse-to-fine tracking strategy with sparse Gaussian selection; and Gaussian-SLAM [20] utilizes DROID-SLAM poses to manage multiple 3D Gaussian submaps.
+These approaches showcase the versatility of 3D Gaussian Splatting in various SLAM settings.
+However, integrating semantics into Gaussian Splatting-based SLAM remains challenging.
+Jointly optimizing geometry, appearance, and semantics involves heterogeneous value ranges and significantly increases memory and computational demands.
+GS3LAM [66] fuses 2D semantic labels into an RGB-D SLAM framework, with embeddings sized according to the number of semantic classes.
+SGS-SLAM [67] appends RGB 3-channels to support semantic visualization.
+SemGauss-SLAM [68] leverages flat semantic embeddings supervised by foundation models, though this introduces high computational cost and overlooks the hierarchical structure of real-world semantics.
+To address these limitations, we propose a compact and efficient solution that fully exploits the hierarchical nature of semantic categories.
+Our method encodes semantic information into a hierarchical symbolic tree, generated through the synergy of large language models and 3D generative priors, and learns this semantic representation end-to-end within the SLAM framework operation.
 III Method
-Figure 2: Left: Overview of the Hier-SLAM++ pipeline.
-The global 3D Gaussian map is initialized using the first frame of the video stream input.
+Figure 2: Overview of the Hier-SLAM++ pipeline.
+Top Left: Hierarchical representation of semantic information.
+The Tree Generation process leverages the capabilities of both LLMs and 3D Generative Models (see Section III-A).
+This hierarchical tree is used to establish a symbolic coding for each Gaussian primitive.
+Additionally, we introduce a novel loss function that combines Intra-level Loss $L_{\text{Intra}}$ and Inter-level Loss $L_{\text{Inter}}$ to optimize the hierarchical semantic representation (see Section III-B).
+Bottom Left: An example of hierarchical semantic rendering.
+Right: The global 3D Gaussian map is initialized using the first frame of the video stream input.
 The system then alternates between the Tracking and Mapping steps as new frames are processed (see Section III-D).
 In the RGB-D setting, depth is directly obtained from sensor input, whereas in the monocular setting the 3D feed-forward method (DUSt3R) is used to generate a geometric prior for depth estimation (see Section III-C).
-Top Right: Hierarchical representation of semantic information.
-The Tree Generation process leverages the capabilities of both LLMs and 3D Generative Models in a top-to-bottom manner.
-This hierarchical tree is used to establish a symbolic coding for each Gaussian primitive (see Section III-A).
-Additionally, we introduce a novel loss function that combines Inter-level Loss $L_{\text{Inter}}$ and Cross-level Loss $L_{\text{Cross}}$ to optimize the hierarchical semantic representation (see Section III-B).
-Bottom Right: An example of hierarchical semantic rendering.
-The entire pipeline of our method is illustrated in Fig.
+The overall pipeline of our method is illustrated in Fig.
 2.
-Within Hier-SLAM++, we model the semantic space as a hierarchical tree, where semantic information is represented as a root-to-leaf hierarchical symbolic structure.
-This representation not only comprehensively captures the hierarchical attributes of semantic information, but also effectively compresses storage requirements, while enhancing semantic understanding with valuable scaling-up capability.
-To construct an effective tree, we leverage both LLMs and 3D Generative Models to integrate semantic and geometric information.
-Next, each 3D Gaussian primitive is augmented with a hierarchical semantic embedding code, which can be learned in an end-to-end manner through the proposed Inter-level and Cross-level semantic losses.
-Furthermore, our method is extended to support monocular input by incorporating geometric priors extracted from the feed-forward method, thereby eliminating the dependency on depth information.
-The following sections provide detailed descriptions for each part of our Hier-SLAM++.
-III-A Hierarchical representation
-We propose hierarchical symbolic representation for semantic information to provide improved efficiency, compression, and compactness for semantic SLAM, making it well-suited for multi-parameter joint optimization and enabling scalability to complex scenes.
-In this section, we provide a detailed explanation of the hierarchical tree parametrization and its generation process.
-III-A1 Tree Parametrization
-The hierarchical tree $\bm{G}$ adopted in Hier-SLAM++ is represented as the set of vertices and edges, $\bm{G}=(\bm{V},\bm{E})$.
-The vertices set $\bm{V}=\cup_{l=0}^{L}\left\{\bm{v}_{l}\right\}$ comes from the classes from all tree levels, where $\left\{\bm{v}_{l}\right\}$ represents the set of nodes at the $l$-th level of the tree, and we use $L$ to represent depth of the whole tree.
-The edge set $\bm{E}=\cup_{m=0}^{L-1}\left\{\bm{e}_{m}\right\}$ captures the subordinate relationships, reflecting semantic attribution, size message, and geometric knowledge.
-For the $i$-th semantic class $g^{i}$, which is treated as a single leaf node in the tree view, its hierarchical expression is:
+At the core of our approach, we model the entire semantic space as a compact hierarchical structure, where each semantic class is represented by a symbolic hierarchy.
+To construct a generative hierarchical tree that captures both semantic relations and associated geometric attributes, we leverage Large Language Models (LLMs) and 3D generative models.
+The hierarchical representation and tree generation process are detailed in Section III-A.
+Based on the proposed structure, semantic information is encoded into the embedding space in two formats: a one-hot representation and a binary representation, both of which are hierarchically organized and compact.
+To support end-to-end training, we introduce hierarchical optimization with two loss functions, Intra-level and Inter-level, which jointly optimize the semantic space based on incoming observations.
+The two proposed tree encoding schemes and hierarchical optimization are discussed in detail in Section III-B.
+In the monocular setting, we leverage geometric priors extracted from a feed-forward 3D model to guide learning during SLAM, incorporating online depth adjustment and geometric supervisory signals, as described in Section III-C.
+Finally, the full semantic Gaussian Splatting SLAM system is presented in Section III-D.
+III-A Hierarchical Representation and Generation
+Leveraging the inherent hierarchy of semantic classes, we propose a hierarchical symbolic representation to encode semantic information in a compact and efficient manner, enhancing the scalability of semantic SLAM system.
+This section presents a detailed description of the proposed hierarchical tree parametrization and its generation process using both LLMs and 3D generative models.
+III-A1 Hierarchical Tree Parametrization
+The proposed hierarchical representation is abstracted as a multi-level tree, formulated as $\bm{G}=(\bm{V},\bm{E})$, where $\bm{V}$ and $\bm{E}$ denote the sets of vertices and edges, respectively, as illustrated in Fig.
+3.
+The vertex set is defined as $\bm{V}=\left\{v_{l}\right\}_{l=0}^{L}$, where $\bm{v}_{l}$ represents the set of semantic classes at level $l$ of the tree hierarchy, spanning from the root level ($l=0$) to the deepest level ($l=L$).
+The edge set $\bm{E}=\left\{e_{m}\right\}_{m=0}^{L-1}$ encodes the subordinate relationships between adjacent levels in the hierarchy, capturing semantic attribution, relative scale, and geometric structure.
+The original semantic classes are represented as leaf nodes located at the deepest level ($l=L$) of the hierarchical tree.
+For the $i$-th semantic class $g^{i}$, its hierarchical representation is defined as:
 $g^{i}=\{v_{l}^{i},e_{m}^{i}\mid l=0,1,...,L;\;m=0,1,...,L-1\},$ (1)
-which corresponds to the root-to-leaf path: $g^{i}=v_{0}^{i}\xrightarrow{e_{0}}v_{1}^{i}\xrightarrow{e_{1}}\cdots\xrightarrow{e_{L-2}}v_{L-1}^{i}\xrightarrow{e_{L-1}}v_{L}^{i}$, as illustrated in Fig.
+which corresponds to a hierarchical path for the semantic meaning: $g^{i}=v_{0}^{i}\xrightarrow{e_{0}}v_{1}^{i}\xrightarrow{e_{1}}\cdots\xrightarrow{e_{L-2}}v_{L-1}^{i}\xrightarrow{e_{L-1}}v_{L}^{i}$.
+For example, the 5-level hierarchical symbolic representation of the class Bed can be expressed as: $g^{\texttt{bed}}=\left\{v_{0}:\texttt{LargeItems}\right\}\rightarrow\left\{v_{1}:\texttt{Furnishings}\right\}\rightarrow\left\{v_{2}:\texttt{BedroomFurniture}\right\}\rightarrow\left\{v_{3}:\texttt{Rectangular}\right\}\rightarrow\left\{v_{4}:\texttt{Bed}\right\}$.
+In this representation, the original semantic label $g^{i}$ is abstracted as a hierarchical path from a coarse category to a fine-grained concept.
+At each level, symbolic nodes $v_{l}^{i}$ represent the semantic attributes of the class, while inter-level relationships, such as inclusion and possession, are described by the edge elements $e_{m}^{i}$.
+As will be demonstrated in Section III-B, with a carefully structured design, the proposed hierarchical representation yields a compact and computationally efficient representation.
+It allows each semantic class to be expressed in a progressive, interpretable, and symbolic manner, incorporating both semantic abstraction and geometric structure.
+Moreover, from a hierarchical perspective, the original flat representation can be regarded as a degenerate case of a single-level tree, enabling a unified formulation.
+For clarity and notational simplicity, we omit the superscripts (i.e., class indices) of variables in the remainder of this section.
+III-A2 Tree Generation via LLMs and 3D Generative Models
+To represent semantic information in a reasonable, comprehensive, and compact manner, the tree structure should integrate both semantic and geometric cues, two fundamental sources of information in unknown environments that are essential for achieving a holistic understanding of the global 3D scene.
+To extract high-level semantic concepts, including functional relationships and abstract descriptors, we employ large language models (LLMs), specifically GPT-4 Turbo [69], which is well known for its strong commonsense reasoning and language understanding abilities.
+On the other hand, to capture general geometric information beyond dataset-specific constraints, we employ 3D generative models, particularly text-to-3D frameworks.
+These models are capable of producing diverse and geometrically accurate 3D objects from textual prompts after appropriate training, offering a unified and generalizable geometric representation.
+Specifically, we adopt MeshGPT [23], a text-to-3D model that learns symbolic 3D embeddings for the local mesh geometry and topology of each object.
+This property makes it particularly well-suited for our goal of constructing geometry-aware hierarchical semantic structures.
+However, the text-to-3D method generates 3D objects within a unified coordinate system but omits explicit size information.
+To address this limitation, we complement the generated geometric shapes with size attributes inferred from large language models (LLMs), enabling a more comprehensive geometric representation for each semantic class.
+Leveraging both large language models (LLMs) and 3D generative models enables the construction of an elaborate and well-structured hierarchical tree, where each level captures a distinct aspect of a class’s attributes—semantic, functional, or geometric, which further enriches the depth and expressiveness of global 3D scene understanding.
+The remainder of this section presents a detailed description of the hierarchical tree generation process.
+The overall process of hierarchical tree generation is illustrated in Fig.
 3.
-Take a semantic class $g=\texttt{Bed}$ as an example, a 5-level tree symbolic coding can be represented as: $\left\{v_{0}:\texttt{LargeItems}\right\}\rightarrow\left\{v_{1}:\texttt{Furnishings}\right\}\rightarrow\left\{v_{2}:\texttt{BedroomFurniture}\right\}\rightarrow\left\{v_{3}:\texttt{Rectangular}\right\}\rightarrow\left\{v_{4}:\texttt{Bed}\right\}$.
-The hierarchical attributes and properties of class $g^{i}$ are represented by the symbolic node information $v_{l}^{i}$, while its relationships such as ‘include’ and ‘possessing’ are represented by the edge information $e_{m}^{i}:\rightarrow$.
-In this way, every semantic class can be coded in a progressive, symbolic hierarchical manner, incorporating both semantic and geometric perspectives.
-Moreover, the standard flat representation can be seen as a single-level tree coding from the hierarchical viewpoint.
-Herein, we shall omit the superscripts (class indexes) of variables for clarity and compact notation.
-III-A2 Tree Generation with LLMs and 3D Generative Models
-To represent semantic information both accurately and comprehensively, an effective tree structure should incorporate information from both semantic and geometric perspectives, as these are the two crucial types of information in an unknown environment, which are essential for achieving a comprehensive understanding of the global 3D environment.
-For capturing semantic meanings, such as functional relationships and summary description labels, we use Large Language Models (LLMs), specifically GPT-4 Turbo [39], known for its strong commonsense reasoning and language understanding capabilities.
-On the other hand, to capture general geometric information without dataset limitations, we employ 3D Generative Models, specifically text-to-3D models, as they have the capability to generate generalized and unified 3D objects with accurate shape information based on text prompts after training.
-We adopt the text-to-3D approach MeshGPT [26], which learns 3D symbolic embeddings with the local mesh geometry and topology of 3D assets, making it well-suited for our purpose.
-However, the text-to-3D method only generates 3D objects within a unified coordinate system, omitting size information.
-To address this issue, we complement the geometric shape information with the size attributes derived from LLMs, to capture the comprehensive geometric information for each semantic class.
-The leverage of both LLMs and 3D Generative Models enables the creation of an elaborate and well-structured hierarchical tree representation, where each level reflects a distinct aspect of a class’s attributes, enhancing the depth and utility of the global 3D understanding.
-The following part provides a detailed explanation of the tree generation process.
-As illustrated in Fig.
-3, the tree generation process follows a top-to-bottom manner.
-At the upper level, the initial grouping divides all classes into several subgroups based on their typical physical sizes in scenes, obtained through prompt queries to the LLM:
+At the top level, the initial grouping divides all semantic classes into several subgroups based on their typical physical sizes in real-world scenes, which are obtained through prompt-based queries to the LLM.
+The generation of level-0 message through this size-based grouping is formulated as:
 $\{v,e\}_{0}=\mathcal{L}^{\text{size}}\{g\}$ (2)
-where $\mathcal{L}^{\text{size}}$ represents the LLMs usage for the size message, and the subscript in $\{v,e\}_{0}$ represents the generation at the initial levels.
-The clustering results categorize classes into groups such as small items, medium items, and large items.
-We provide details of the specific prompt used for the LLM in Appendix 1.
-Next, the classes within each group are used as input prompts to LLMs for further clustering based on their semantic relationships, particularly their functional properties:
+where $\mathcal{L}^{\text{size}}$ denotes the LLM-driven grouping operator.
+The resulting clusters categorize the semantic classes into groups such as SmallItems, MediumItems, and LargeItems.
+Details of the specific prompt used to query the LLM are provided in Appendix-I.
+Next, the semantic classes within each size-based group are used as input prompts to the LLM for further clustering based on semantic relationships, with an emphasis on functional properties.
+This step generates the following levels structure, formulated as:
 $\{v,e\}_{1}=\mathcal{L}^{\text{func}}\{g\}$ (3)
-where $\mathcal{L}^{\text{func}}$ denotes the LLMs clustering based on functional relationships, and the subscript of $\{v,e\}_{1}$ stands for the following levels generation.
-We demonstrate the detail prompt used for the LLM in Appendix 1.
-After multiple levels generation using LLMs, 3D Generative Models are employed to cluster classes based on their geometric shapes.
-Specifically, we adopt MeshGPT [26] to capture the geometric information for each semantic class.
-MeshGPT [26] is a text-to-3D model that encodes object shape and internal topology relationships into latent quantized embeddings, forming a symbolic 3D representation in the shape space.
-Given a text prompt (i.e., the class name) as input to the pre-trained MeshGPT, we extract the latent quantized embeddings to represent the shape information for the class:
-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{z}=\mathcal{E}\{g\}\in\mathbb{N}^{F\times 2}}}$ (4)
-where $g$ is the class name provided as input.
-$\mathcal{E}$ represents the pre-trained MeshGPT encoder, and $\bm{z}$ is the quantized geometric embedding extracted from the pre-trained encoder.
-For each object’s embedding, $F$ is the number of triangle faces used in the mesh representation.
-Since the number of triangle faces varies across different semantic objects, i.e., different values of $F$, we first compute the average shape embedding $\bm{z}_{\text{avg}}\in\mathbb{R}^{1\times 2}$ for each object.
-The averaged embeddings from all classes are then concatenated and subsequently normalized to a standard distribution (zero mean and unit variance) to ensure consistency across classes:
+where $\mathcal{L}^{\text{func}}$ denotes the LLM-driven clustering operator based on functional similarity.
+It is worth noting that $\{v,e\}_{1}$ does not refer strictly to a single hierarchical level.
+Rather, it denotes a generic stage of functional grouping that follows the initial size-based partitioning.
+Depending on the balance and structural design of the overall tree, the functional clustering process may be applied multiple times to generate several intermediate levels.
+This hierarchical refinement enables the representation to achieve maximum semantic compactness while maintaining accuracy and interpretability.
+The specific prompt used for this functional clustering is provided in Appendix-I.
+After generating the preceding semantic substructures using LLMs across multiple levels, 3D generative models are employed to further cluster semantic classes based on their geometric characteristics.
+This geometry-based symbolic representation complements the language-derived hierarchy, enabling a unified structure that incorporates both semantic and geometric information.
+Specifically, we adopt MeshGPT [23] to capture geometric information for each semantic class.
+This text-to-3D framework encodes object shapes and internal topological relationships into latent, quantized embeddings, resulting in a symbolic 3D representation within the geometric shape space.
+Given a text prompt (i.e., the class name) as input to the pre-trained MeshGPT model, we extract the latent quantized embeddings to represent the geometric information of the class:
+$\bm{z}=\mathcal{E}\{g\}\in\mathbb{N}^{F\times 2}$ (4)
+where $\mathcal{E}$ denotes the pre-trained MeshGPT encoder, and $\bm{z}$ is the quantized geometric embedding extracted for class $g$.
+Here, $F$ represents the number of triangle faces used in the mesh representation [23].
+Since the number of triangle faces $F$ varies across semantic objects, we first compute an average shape embedding $\bm{z}_{\text{avg}}\in\mathbb{R}^{1\times 2}$ for each class.
+These averaged embeddings are then concatenated and normalized to a standard distribution (i.e., zero mean and unit variance) to ensure consistency across classes:
 $\tilde{\bm{z}}=\mathcal{N}\{\mathcal{C}(\bm{z}_{\text{avg}})\}$ (5)
 where $\mathcal{N}$ and $\mathcal{C}$ denote the normalization and concatenation operations, respectively.
-Next, the normalized embeddings $\tilde{\bm{z}}$ are clustered using K-means++, denoted by the operator $\mathcal{K}$.
-To address the lack of descriptive labels for each clustering result, we further employ LLMs to summarize and refine the clustering outcomes:
+The resulting normalized embeddings $\tilde{\bm{z}}$ are subsequently clustered using the K-means++ algorithm [70], denoted by the operator $\mathcal{K}$.
+To address the lack of descriptive labels for the resulting clusters, we further employ an LLM-based summarization module to interpret and refine the clustering outcomes:
 $\{v,e\}_{2}=\mathcal{L}^{\text{sum}}\{\mathcal{K}(\tilde{\bm{z}})\}$ (6)
-where $\mathcal{L}^{\text{sum}}$ represents the use of LLMs for summarization.
-The generated descriptive labels, such as ”box”, ”flat”, ”soft”, summarize the common geometric attributes within each grouping result, further enhancing the interpretability of the hierarchical tree.
-The details are illustrated in Appendix 1.
-Whenever LLMs are employed, we ensure balanced, meaningful, and descriptive outputs through carefully designed prompt inputs, as detailed in Appendix 1.
-Furthermore, an additional LLM validator [25] is utilized after each generation level to verify clustering completeness and accuracy, ensuring that no classes are missing or incorrectly included (see Appendix 1 for detail).
-Finally, a global tree evaluation is conducted to detect and resolve any duplicates or omissions across tree levels, ensuring a fully balanced and complete hierarchical tree structure.
-We use manual inspection as the final verification to ensure the correctness and rationality of the holistically generated tree.
+where $\mathcal{L}^{\text{sum}}$ denotes the use of an LLM to generate semantic summaries for each geometric cluster.
+The resulting descriptive labels, such as “box”, “flat”, or “soft”, capture the dominant geometric attributes within each group, thereby enhancing the interpretability of the hierarchical structure.
+Details of the LLM prompts and summarization process are provided in Appendix-I.
+Whenever LLMs are employed, we ensure that the outputs are balanced, meaningful, and descriptively coherent through carefully designed prompt inputs, as detailed in Appendix-I.
+To further ensure the structural correctness within each level of the hierarchical tree, we introduce a loop-based critic operation performs multiple iterations.
+In each iteration, an LLM is followed by a validator module that checks the quality of the generated clusters.
+Specifically, within each LLM-based layer, the validator assesses the clustering results in terms of completeness and semantic consistency by comparing them against the input prompt.
+Correctly grouped nodes are retained, while erroneously included classes are removed.
+Any omitted classes are compiled into a new prompt and fed into the next iteration of LLM-based clustering.
+This process repeats until no omitted classes remain at the current level, ensuring a complete and semantically accurate hierarchical structure.
+Finally, a comprehensive validation process is applied to the entire hierarchical structure, including an automated check to detect and resolve duplicates or omissions across levels, followed by a manual inspection to ensure correctness and conceptual consistency.
+This ensures a fully balanced and semantically coherent tree.
 Figure 3: Visualization of the hierarchical tree structure and semantic embedding.
@@
 For tree generation, we utilize both LLMs and 3D Generative Models to extract and group messages.
-For hierarchical semantic embedding, we represent semantic information using a root-to-leaf symbolic path.
-We propose two types of representations: the one-hot representation and the binary representation, which can reduce the original dimensionality by up to $O(\log^{2}N)$.
-III-B Hierarchical optimization
-Based on the generated hierarchical tree structure, the semantic information is represented as compact hierarchical symbolic nodes, and further integrated into the 3D Gaussian primitives, learned in an end-to-end manner with the input frame stream.
-To ensure accurate semantic understanding, we introduce a semantic hierarchical loss, which consists of Inter-level and Cross-level losses.
-These losses are incorporated into the optimization process during the online SLAM operation.
-The following sections provide a detailed explanation of the proposed tree encoding and the loss calculation.
-III-B1 Tree Encoding
-As introduced in Eq.
-(1), the semantic class in Hier-SLAM++ is represented hierarchically in a root-to-leaf symbolic manner based on the generated tree.
-This effectively compacts semantic information, facilitates SLAM joint optimization, and provides valuable scalability for semantic understanding in unknown environments.
-Specifically, we propose two types of hierarchical semantic embedding approaches, both of which compact the original semantic information.
-One-hot representation: As demonstrated in Fig. 3, the hierarchical semantic embedding $\bm{h}$ is composed of the one-hot embeddings across all levels:
+For hierarchical semantic embedding, we represent semantic information using a hierarchical symbolic path.
+We propose two types of representations: the one-hot representation and the binary representation, which can reduce the original dimensionality by up to $O(\log N)$.
+III-B Hierarchical Semantic Coding and Optimization
+Based on the generated hierarchical tree structure, semantic information is encoded as compact, symbolic hierarchical codes and integrated into the 3D Gaussian primitives, which are learned in an end-to-end fashion from the input image stream.
+To ensure accurate and structured semantic understanding, we introduce a semantic hierarchical loss, composed of both Intra-level and Inter-level components.
+These losses are incorporated into the overall optimization during the online SLAM process.
+The following sections provide a detailed description of the hierarchical semantic encoding and the proposed optimization framework.
+III-B1 Hierarchical Semantic Coding
+As introduced in Section III-A, we model the semantic space as a hierarchical tree structure.
+Within this hierarchical space, each semantic class is represented in a structured and compact form that preserves essential semantic information.
+The proposed hierarchical semantic coding enhances the joint optimization of the SLAM system and offers valuable scalability for semantic understanding in the complex environments.
+Specifically, we introduce two types of hierarchical semantic coding strategies, both of which compress the original flat semantic labels while retaining crucial semantic messages.
+One-hot representation: As illustrated in Fig. 3, the hierarchical semantic embedding $\bm{h}$ is constructed by concatenating one-hot embeddings from all levels of the tree:
 $\bm{h}=\mathcal{C}(\bm{h}_{l})\in\mathbb{B}^{N},\quad\bm{h}_{l}\in\mathbb{B}^{n}$ (7)
-where $l$ represents the $l$-th level of the tree, and $\mathcal{C}$ denotes the concatenation operation.
-For each level’s representation $\bm{h}_{l}$, we use $n$-dimensional one-hot codes (in colored boxes) to encode the symbolic nodes (in colored circles) from the tree of this level, as shown in Fig.
+where subscript denotes the $l$-th tree level, and $\mathcal{C}$ represents the concatenation operator.
+At each level, $\bm{h}_{l}$ is an $n$-dimensional one-hot vector (as shown in colored boxes in Fig.
+3) used to represent symbolic nodes (colored circles) within that level.
+The dimension $n$ corresponds to the maximum number of nodes at level $l$, indexing node representations from 00 to $n-1$.
+The total dimensionality of the hierarchical semantic embedding is given by $N=\sum n$, i.e., the sum over all levels.
+Compared to the original flat representation of dimension $N_{\text{original}}$, this hierarchical structure enables an information-preserving reduction of up to $O(\log N)$ in representational complexity.
+The resulting compact semantic code significantly reduces memory usage and facilitates efficient joint optimization in the SLAM system.
+Binary representation: To further improve compactness, we propose a binary version of the hierarchical semantic representation.
+The binary embedding $\bm{b}$ is constructed by concatenating the binary vectors $\bm{b}_{l}$ across all levels of the hierarchical tree:
+$\bm{b}=\mathcal{C}(\bm{b}_{l})\in\mathbb{B}^{K},\quad\bm{b}_{l}\in\mathbb{B}^{k},\quad\text{where }k=\left\lceil\log_{2}n\right\rceil$ (8)
+where $\bm{b}_{l}$ is the binary code converted from the one-hot embedding $\bm{h}_{l}$ at level $l$, i.e., $\bm{b}_{l}=\text{Binary}(\bm{h}_{l})$, as illustrated in the final row of Fig.
 3.
-The dimension $n$ corresponds to the maximum number of nodes at the $l$-th level.
-Using the one-hot representation, the overall dimension of the semantic embedding is the sum of the dimensions across all levels, given by $N=\sum n$.
-This enable a maximum reduction of up to $O(\log N)$ compared to the original dimension $N_{\text{original}}$ in the flat representation.
-This hierarchical representation effectively compacts the semantic embedding, reducing storage requirements and facilitating joint optimization.
-Binary representation: We propose a further compact version of our hierarchical semantic representation in this section.
-The binary representation of the hierarchical semantic embedding $\bm{b}$ consists of the binary embeddings $\bm{b}_{l}$ across all tree levels:
-$\bm{b}=\mathcal{C}(\bm{b}_{l})\in\mathbb{B}^{K},\quad\bm{b}_{l}\in\mathbb{B}^{k}$ (8)
-where $\bm{b}_{l}$ is the binary representation converted from the one-hot semantic embedding $\bm{h}_{l}$ at each level: $\bm{h}_{l}\rightarrow\bm{b}_{l}$, as depicted in the last row example in Fig.
-3.
-For instance, if the one-hot embedding has a dimension of $n=8$, this corresponds to a maximum of $8$ unique codes at this level (covering values from 00 to $7$).
-After binarization, the dimension of the compact version is reduced to $k=3$, since $82^{3}=$ 8.
-By leveraging binary representation, the hierarchical coding dimension at each level is reduced to $k=\log n$, compared to the one-hot encoding with dimension $n$.
-Consequently, the total semantic encoding dimension across all levels is given by $K=\sum k$, resulting in a more compact representation.
+For example, if the one-hot vector has dimension $n=8$ (representing $8$ distinct semantic nodes indexed from 00 to $7$), its binary representation requires only $k=3$ bits for $82^{3}=$ 8.
+More generally, the binary code length at each level is reduced to $k=\lceil\log_{2}n\rceil$, compared to the $n$-dimensional one-hot encoding.
 III-B2 Loss Calculation
-To fully optimize the hierarchical semantic message both within each level and across different levels of the tree, the proposed hierarchical loss consists of Inter-level loss $L_{\text{Inter}}$ and Cross-level loss $L_{\text{Cross}}$:
-$L_{\text{{Semantic}}}=\omega_{1}L_{\text{{Inter}}}+\omega_{2}L_{\text{{Cross}}}$ (9)
+To effectively and efficiently optimize the proposed hierarchical semantic codings, we introduce a hierarchical loss formulation that supervises semantic embedding both within each level (intra-level) and across levels (inter-level).
+Specifically, the total semantic loss is defined as:
+$L_{\text{{Semantic}}}=\omega_{1}L_{\text{{Intra}}}+\omega_{2}L_{\text{{Inter}}}$ (9)
 where $\omega_{1}$ and $\omega_{2}$ are balancing coefficients for each loss term.
-Specifically, the Inter-level loss $L_{\text{Inter}}$ is employed within each level’s embedding $\bm{h}^{l}$ to ensure the optimization of semantic information within each level:
-$L_{\text{{Inter}}}=\sum_{l=0}^{L}L_{\text{{ce}}}(\mathcal{S}(\bm{h}^{l}),\mathcal{P}^{l})$ (10)
+Specifically, the intra-level loss $L_{\text{Intra}}$ is employed within each level’s embedding $\bm{h}^{l}$ to ensure the optimization of semantic information within each level:
+$L_{\text{{Intra}}}=\sum_{l=0}^{L}L_{\text{{ce}}}(\mathcal{S}(\bm{h}^{l}),\mathcal{P}^{l})$ (10)
 where $L_{\text{{ce}}}$ indicates the cross-entropy semantic loss, and $\mathcal{P}^{l}$ denotes the semantic ground truth at level $l$.
 $\mathcal{S}$ stands for the Softmax operation, converting embeddings into probabilities.
-For our proposed compact version with binary representation $\bm{b}^{l}$, we replace the original cross-entropy loss with binary cross-entropy loss $L_{\text{bce}}$ for the binary Inter-level loss calculation:
+For our proposed compact version with binary representation $\bm{b}^{l}$, we employ the binary cross-entropy loss $L_{\text{bce}}$, instead of the standard cross-entropy, for computing the binary intra-level loss:
 $L_{\text{Inter}}^{\text{bin}}=\sum_{l=0}^{L}L_{\text{{bce}}}(\mathcal{S}(\bm{b}^{l}),\mathcal{P}^{l})$ (11)
-To enforce accurate understanding across tree levels, we introduce the Cross-level loss, which is computed using a semantic decoder based on the holistic hierarchical codes $\bm{h}$.
-Specifically, we first perform a 2D convolution operation on the semantic embedding, following an activation layer (i.e., ReLU).
-Then, a second 2D convolution layer is applied to map the hidden embedding into the flat coding, followed by a softmax operation to convert the embeddings into probabilities.
-The overall Cross-level loss $L_{\text{Cross}}$ is defined as:
-$L_{\text{Cross}}=L_{\text{ce}}\big{(}\mathcal{S}(\text{decoder}(\bm{h})),\mathcal{P}\big{)},$ (12)
-where $\mathcal{P}$ represents the semantic ground truth, and the definition of $\text{decoder}(\bm{h})$ is:
+To ensure accurate semantic understanding across all levels of the hierarchical structure, we introduce the inter-level loss $L_{\text{Inter}}$, which enforces consistency between the hierarchical representation and the original flat semantic distribution.
+To this end, we employ a lightweight decoder to transform the holistic hierarchical embedding $\bm{h}$ into flat semantic predictions.
+Specifically, the decoder consists of two stacked 2D convolutional layers, with a ReLU activation applied between them.
+The output is then passed through a softmax function to produce a probability distribution over the flat semantic classes:
 $\text{decoder}(\bm{h})=\text{Conv}\big{(}\text{ReLU}(\text{Conv}(\bm{h}))\big{)}$ (12)
+The inter-level loss is defined as:
+$L_{\text{Inter}}=L_{\text{ce}}\big{(}\mathcal{S}(\text{decoder}(\bm{h})),\mathcal{P}\big{)},$ (13)
+where $\mathcal{S}(\cdot)$ denotes the softmax function, $\mathcal{P}$ is the ground truth flat semantic label, and $L_{\text{CE}}$ is the standard cross-entropy loss.
 III-C Monocular setting with geometric priority
-Accurate depth information provides valuable geometric cues for global map reconstruction and pose estimation in a SLAM system.
-In an RGB-D setting, depth signals are directly obtained from sensors.
-However, in a monocular setting without depth information, SLAM [28, 29, 30, 31] typically relies on triangulation to initialize the global mapping.
-Yet, due to the inherent coupling between pose estimation and mapping, this approach often suffers from relatively lower accuracy compared with RGB-D based methods.
-In our Hier-SLAM++, we employ a feed-forward method as a geometric prior.
-Specifically, we leverage DUSt3R [27] to provide depth geometric priors for monocular SLAM.
-DUSt3R is able to generate both a 3D point map and pose estimations from several sparse-view image inputs, which can serve as the geometric prior for the monocular SLAM.
-The following section provides details on the monocular setting in our Hier-SLAM++.
+Accurate depth information provides valuable geometric cues for global map reconstruction and pose estimation in SLAM systems.
+While RGB-D methods directly acquire depth from sensors, monocular SLAM approaches [28, 29, 30, 31] must rely on multi-view triangulation to infer depth and initialize mapping, which increases the difficulty of simultaneously performing localization and mapping.
+To address this limitation, a number of methods [71, 72, 73, 74, 75, 76] introduce depth priors from learned depth networks into SLAM systems, achieving promising results.
+More recently, a new trend has emerged: 3D feed-forward models, which can generate accurate global 3D representations of previously unseen environments after training.
+In our work, we adopt a precise and highly generalizable 3D feed-forward method to obtain the geometric prior used in monocular SLAM.
+Specifically, we employ DUSt3R [24], which produces both 3D point clouds and pose estimations from sparse-view image inputs, serving as a strong geometric prior in our system.
+Details of the monocular SLAM setting in Hier-SLAM++ are provided in the following section.
 III-C1 Geometric prior acquisition
@@
 $\{I_{i}\;|\;i=i_{0}+\delta\cdot(s-1),\;s=1,\dots,S\}$ (14)
-Each image set is then fed into DUSt3R to generate geometric information for every image frame, specifically the depth maps.
-Although DUSt3R is also able to provide pose predictions for each input frame, we observe that the accuracy of estimated pose is relatively low—performing even worse than a constant velocity pose prediction in the SLAM system based on our experiments.
-Therefore, only the estimated geometric information from DUSt3R is leveraged in our method.
+Each image set is then fed into DUSt3R to generate geometric information, specifically the depth maps.
+While DUSt3R is also capable of estimating camera poses for each input frame, our experiments reveal that its pose predictions exhibit relatively low accuracy, often performing worse than a simple constant-velocity pose model when integrated into a SLAM pipeline.
+Therefore, we retain only the geometric (depth) information predicted by DUSt3R, as it provides more reliable guidance for our SLAM system.
 III-C2 Geometric prior correction
-The scale of the estimation for each image set varies and also differs from the scale of the online global map estimation during SLAM operation, requiring an additional geometric prior correction operator.
-Firstly, we calculate the initial scale and shift parameters by comparing the rendering results from the current global map prediction with the depth prior obtained from DUSt3R.
-Specifically, the depth map $D_{c}$ and the silhouette map $S_{c}$ for current frame $I_{c}$ are rendered based on the global Gaussian map $G$ as well as the pose estimation $\mathbf{T}_{c}$ for the current frame:
+The depth predicted by DUSt3R is scale-ambiguous and lacks metric consistency across image sets.
+This leads to differences in scale between the predicted geometry and the SLAM system’s estimated global map, which evolves over time.
+We therefore incorporate a geometric prior correction step to resolve this scale inconsistency.
+To align the scale of the depth prior with the current global map, we first estimate the scale and shift parameters by comparing the rendered depth from the SLAM map with the predicted depth from DUSt3R.
+Specifically, for each input frame $I_{c}$, we render a depth map $D_{c}$ and silhouette map $S_{c}$ from the global 3D Gaussian map $G$ using the current camera pose $\mathbf{T}_{c}$:
 $\{D_{c},S_{c}\}=\mathcal{R}(G,\mathbf{T}_{c})$ (15)
 where $\mathcal{R}$ represents the rasterization procedure for the 3D Gaussian map.
-Next, the global scale and shift parameters $\lambda_{c},\tau_{c}$ are calculated by comparing the predicted depth $\hat{D}_{c}$ from DUSt3R with the rendered depth $D_{c}$ using the least squares estimation on the visible regions $M_{c}$ defined by $S_{c}$:
+We then compute the global scale and shift parameters $\lambda_{c}$ and $\tau_{c}$ by aligning the DUSt3R-predicted depth $\hat{D}c$ to the rendered depth $D_{c}$ using least squares estimation over the visible regions $M_{c}$, as defined by the silhouette map $S_{c}$:
 $\{\lambda_{c},\tau_{c}\}=\arg\min\left(\hat{D}_{c}^{\text{align}}-D_{c}\right)^{2}M_{c}$ (16)
-where $\hat{D}_{c}^{\text{align}}=\lambda_{c}\cdot\hat{D}_{c}+\tau_{c}$.
-Secondly, the estimated parameters are fed into the optimization as initialization and further optimized jointly within the map optimization via the depth supervision loss:
-$\{\lambda_{c},\tau_{c},G\}=\arg\min|\hat{D}_{c}^{\text{align}}-D_{c}|M_{c}$ (17)
-In this way, the depth prediction can be continuously refined online with increasing observations, providing geometric supervisory signals during SLAM optimization.
-This helps guide global reconstruction and pose estimation, leading to improved results throughout the SLAM process.
+where $\hat{D}_{c}^{\text{align}}=\lambda_{c}\cdot\hat{D}_{c}+\tau_{c}$ denotes the aligned depth.
+This simple affine transformation is sufficient to resolve the scale ambiguity inherent in the geometric priors obtained from DUSt3R, while preserving their relative 3D structure for integration into the monocular SLAM pipeline.
+Secondly, the calculated scale and shift parameters are used as initial values and are jointly optimized with the 3D Gaussians during the map optimization process, guided by a depth supervision loss:
+$\{\lambda_{c},\tau_{c},G\}=\arg\min\big{(}|\hat{D}_{c}^{\text{align}}-D_{c}|M_{c}+L_{\text{reg}}\big{)}$ (17)
+where $L_{\text{reg}}$ is a regularization term defined as the absolute deviation between the optimized parameters and their initialization, preventing overfitting or convergence to poor local minima.
+By incorporating this optimization into the SLAM operation, the depth prior can be continuously refined online as new observations arrive, effectively providing geometric supervision throughout the SLAM process.
+This not only improves the quality of the global map reconstruction but also enhances the accuracy of camera pose estimation.
 III-D Semantic Gaussian Splatting SLAM
-The global semantic 3D Gaussian Splatting SLAM system is illustrated in the left part of Fig.
+The global semantic 3D Gaussian Splatting SLAM system is illustrated in Fig.
 2.
-Based on the established semantic 3D Gaussian representation, global 3D reconstruction (Mapping) and pose estimation (Tracking) are alternately performed with the video stream input.
-For the RGB-D setting, depth information for each image frame is obtained directly from depth sensors.
-For the monocular setting, we employ a feed-forward method to obtain the geometric prior, as introduced in the previous section.
-Further details of our Hier-SLAM++ system are provided in the following section.
+Built upon the semantic 3D Gaussian representation, global reconstruction (Mapping) and camera pose estimation (Tracking) are alternately performed over the input video stream.
+In the RGB-D setting, depth information for each frame is directly acquired from the depth sensor.
+In the monocular setting, a feed-forward model is employed to generate geometric priors, as described in the previous section.
+The construction of the semantic 3D Gaussians and the overall semantic SLAM system are detailed in the following sections.
 III-D1 Semantic 3D Gaussian representation
-The semantic 3D Gaussian representation is represented as the Gaussian primitives with the hierarchical semantic embedding.
-Motivated by [16], we adopt the isotropic and view-independent Gaussian integrated with the semantic embeddings.
-Each semantic Gaussian is composed of color $\bm{c}$, the center position $\bm{\mu}$, the radius $r$, the opacity $o$, and its semantic embedding $\bm{h}$.
-And the influence of each Gaussian $G(\bm{X})$ [13] according to the standard Gaussian equation is:
+Our semantic 3D Gaussian representation encodes both geometric and semantic information using Gaussian primitives augmented with hierarchical semantic embeddings.
+Motivated by [16], we adopt isotropic, view-independent Gaussians and integrate them with semantic feature vectors.
+Each semantic Gaussian is parameterized by its color $\bm{c}$, center position $\bm{\mu}$, radius $r$, opacity $o$, and semantic embedding $\bm{h}$.
+The influence of each Gaussian $G(\bm{X})$ [13] at 3D point $\bm{X}$ follows the standard Gaussian formulation:
 $G(\bm{X})=o\>\exp\left(-\frac{||\bm{X}-\bm{\mu}||^{2}}{2r^{2}}\right)$ (18)
 where $\bm{X}$ stands for the 3D point.
-Following [13], every semantic 3D Gaussian primitive within the current global map is projected to the 2D image space using the tile-based differentiable $\alpha$-compositing rendering.
+Following [13], every semantic 3D Gaussian primitive within the current global map is projected onto the image plane using the tile-based differentiable $\alpha$-compositing rendering.
 The semantic map $H$ is rendered as follows:
-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{H=\sum_{i=1}^{n}\bm{h}_{i}\alpha_{i}(\bm{X})T_{i}\quad\text{with}\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{X}))}}$ (19)
+$H=\sum_{i=1}^{n}\bm{h}_{i}\alpha_{i}(\bm{X})T_{i}\quad\text{with}\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{X}))$ (19)
 The rasterized color image $C$, depth image $D$, and the silhouette image $S$ are defined as follows:
 $C=\sum_{i=1}^{n}\bm{c}_{i}\alpha_{i}(\bm{X})T_{i},\;D=\sum_{i=1}^{n}\bm{d}_{i}\alpha_{i}(\bm{X})T_{i},\;S=\sum_{i=1}^{n}\alpha_{i}(\bm{X})T_{i}$ (20)
-Different from previous work [16], which uses separate forward and backward Gaussian modules for different parameters, our Hier-SLAM++ system adopt a unified forward and backward modules that processes all parameters, including semantics, color, depth, and silhouette images—significantly improving overall runtime efficiency.
+Different from previous work [16], which uses separate forward and backward Gaussian modules for different parameters, our Hier-SLAM++ system adopts a unified forward and backward modules that processes all parameters, including semantics, color, depth, and silhouette images, which significantly improving overall runtime efficiency.
 III-D2 Tracking
@@
 Following that, a pose optimization is performed with the global Gaussian map $G$ fixed:
-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathbf{T}_{c}=\arg\min_{\mathbf{T}_{c}}L_{\text{Track}}(G,\mathbf{T}_{c})}}$ (21)
+$\mathbf{T}_{c}=\arg\min_{\mathbf{T}_{c}}L_{\text{Track}}(G,\mathbf{T}_{c})$ (21)
 where the tracking loss includes the rendering color and depth losses:
@@
 III-D3 Mapping
-The global map information, including semantic information, is optimized in the mapping procedure while keeping camera poses fixed:
-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{G=\arg\min_{G}L_{\text{Map}}(G,\mathbf{T}_{i})}}$ (23)
+The global map information, including semantic information, is optimized in the mapping procedure while keeping camera poses $\mathbf{T}_{i}$ fixed:
+$G=\arg\min_{G}L_{\text{Map}}(G,\mathbf{T}_{i})$ (23)
 The mapping optimization losses include depth, color, and semantic losses:
@@
 Additionally, we compare with state-of-the-art semantic SLAM methods, covering both NeRF-based and Gaussian-based techniques, to highlight its capability in hierarchical semantic representation and scalability.
-The experiments are conducted in the Nvidia L40S GPU.
+Experiments are conducted on an NVIDIA L40S GPU, while runtime tests are evaluated on both the L40S and RTX 4090 to enable broader comparisons.
 For experimental settings, the semantic embedding of each Gaussian primitive is initialized randomly.
@@
 For mapping, the loss weights are set as follows: $w_{3}=1.0$, $w_{4}=0.5$, and $w_{5}=0.2$.
+The two types of semantic embeddings proposed in our method, one-hot and binary, are evaluated and reported across multiple experiments, including tracking, mapping, rendering, runtime, and semantic understanding, providing a comprehensive ablation study on the effectiveness and compactness of the hierarchical semantic representation.
 TABLE I: RGB-D Tracking performance ATE RMSE (cm) on the Replica.
-Best results are highlighted as FIRST, SECOND.
+Best results are highlighted as first, second, third.
 Methods Avg. R0 R1 R2 Of0 Of1 Of2 Of3 Of4
@@
 III.
-Since the TUM-RGBD dataset does not contain semantic ground truth, we evaluate our method based on tracking accuracy here.
-Despite the low-quality RGB and depth inputs, as well as the absence of semantic information, all of which could negatively impact overall performance, our method still achieves similar performance to the state-of-the-art approach [16], with a localization error of less than $0.2$ cm, while operating at a significantly faster runtime.
-A detailed analysis of the runtime performance will be provided in the following section.
+As this dataset does not provide semantic ground truth, we evaluate our method solely based on tracking accuracy.
+Without access to semantic information, our system operates as a standard visual SLAM pipeline.
+As a result, the tracking accuracy becomes comparable to the state-of-the-art method [16] in the TUM-RGBD dataset.
 TABLE II: RGB-D Tracking performance ATE RMSE (cm) on the Scannet.
@@
 ElasticFusion [83] 6.91 2.53 6.83 21.49 1.17 2.52
-ORB-SLAM2 [29] 1.98 1.60 2.20 4.70 0.40 1.00
+ORB-SLAM3 [30] 2.21 1.60 2.37 5.83 0.32 0.94
 NICE-SLAM [59] 15.87 4.26 4.99 34.49 31.73 3.87
@@
 […]
-IV-C Global Mapping Performance
-Following previous methods [25, 23, 24], we evaluate the depth error on the Replica dataset [41] to assess overall mapping performance, as it provides the highest-quality depth maps for comparison.
-With an accurate understanding of semantics, our method effectively learns precise global geometry in unknown environments, benefiting both tracking and mapping in the semantic SLAM system.
-Our method achieves superior results in 7 out of 8 sequences and maintains a low average depth error across all sequences, with a difference of approximately $0.1$ cm compared to the best performing.
-These results demonstrate that our approach have strong semantic mapping capability, highlighting our superior mapping performance.
-IV-D Rendering Quality
-We present the quantitative analysis of the rendering performance of our proposed method in Tab.
-V.
-Following previous works, Point-SLAM [45] and NICE-SLAM [33], the evaluation is conducted on the input views.
-As shown in Tab.
-V, our method achieves superior image rendering performance compared to state-of-the-art approaches across all 8 sequences, as well as in the overall average evaluation.
-As discussed in previous sections, incorporating semantic information introduces a significant computational burden to the joint optimization process within the SLAM system.
-However, our method demonstrates notably improved rendering performance compared to all existing methods, including the non-semantic SLAM approaches.
-Although the additional computation for semantic learning increases the optimization complexity, the accurate semantic understanding enhances both geometric and appearance learning through joint optimization, effectively improving rendering performance.
-The visual comparison of rendered images is presented in Fig.
-6.
-Benefiting from precise semantic information learning and the improved SLAM system, the rendering results exhibit significantly improved detail preservation, such as more accurate and well-defined lamp boundaries in the first and fifth columns, and precise details even in fine structures like shutters in the second column.
-Moreover, our method produces high-quality renderings with fewer noise artifacts, as evidenced by the improved rendering quality of the floor in the fourth column and the chair in the sixth column.
-This demonstrates that better mapping performance leads to enhanced rendering quality.
-The joint optimization of semantic losses and photometric loss is mutually beneficial, facilitating more accurate global 3D reconstruction through semantic learning, which in turn improves the rendering performance of our Hier-SLAM++.
-TABLE IV: Reconstruction performance Depth L1 (cm) on Replica. Best results are highlighted as first, second.
+TABLE IV: Reconstruction performance Depth L1 (cm) on Replica. Best results are highlighted as first, second, third.
 Methods Avg. R0 R1 R2 Of0 Of1 Of2 Of3 Of4
@@
 Hier-SLAM++ (binary) 0.49 0.59 0.40 0.40 0.30 0.17 0.52 0.95 0.57
-IV-E Running time
-The runtime results for the compared methods are provided in Tab.
-VI, covering both NeRF-based and Gaussian Splatting-based SLAM methods.
-Since the compared baseline using a flat representation (Hier-SLAM++-flat) cannot run successfully on an NVIDIA RTX 4090 due to GPU memory limitations, we report the runtime separately for RTX 4090 and L40S.
-In the first block of Tab.
-VI, we observe that, compared to NeRF-based methods [33], Gaussian-based state-of-the-art methods [16] achieve significantly higher operating speeds, benefiting from the fast rendering and efficient optimization capabilities inherent to the new 3D representation, i.e., Gaussian Splatting.
-Furthermore, with the carefully optimized forward and backward Gaussian rendering module introduced in our system, which serves as the core component of Gaussian Splatting-based SLAM systems, our method (Hier-SLAM++- w/o sem) achieves up to 2.4× faster tracking and 2.2× faster mapping compared to the best-performing method [16].
-In the second block of Tab.
-VI, when integrating semantic information, our method maintains high efficiency by leveraging hierarchical semantic coding (Hier-SLAM++-one-hot), achieving nearly 3× faster tracking compared to flat semantic coding (Hier-SLAM++-flat).
-Compared to the state-of-the-art method [25], our Hier-SLAM++ attains similar operational speed but significantly outperforms in semantic segmentation accuracy.
-Thanks to the improved semantic decoder in our hierarchical semantic loss, our method achieves a $13\%$ mIoU improvement while incurring only an additional $0.04$ s per frame for tracking and $2$ s per frame for mapping.
-Further details on this improvement will be provided in the next section.
-Additionally, our binary version (Hier-SLAM++-binary) is able to achieve more efficient operation performance.
-Notably, our Hier-SLAM++ achieves a rendering speed of 2000 FPS, which increases to 3000 FPS when semantic information is not included.
-TABLE V: Rendering performance PSNR, SSIM, LPIPS on Replica. Best results are highlighted as first, second.
+IV-C Global Mapping Performance
+Following prior work [25, 67, 68], we evaluate depth accuracy on the Replica dataset [78], which provides high-quality ground truth depth maps for reliable evaluation of overall mapping performance.
+The quantitative results are presented in Tab.
+IV.
+Our method achieves the lowest depth error in 7 out of 8 sequences and maintains a consistently low average error across all sequences, with only a $0.1$ cm difference compared to the best-performing baseline.
+By leveraging semantic understanding, our approach enables the learning of precise geometry in unseen environments, which in turn benefits both tracking and mapping within the whole semantic SLAM system.
+These results confirm the strong semantic mapping capability of our method and highlight its superior performance in dense 3D reconstruction.
+IV-D Rendering Quality
+Rendering performance is becoming an increasingly important evaluation metric, particularly for SLAM systems built upon new scene representations, such as neural implicit representations and gaussian splatting.
+Following prior works [62, 59], the evaluation is conducted on the input views of the Replica dataset.
+We present a quantitative analysis of the rendering performance of our proposed method in Tab.
+V.
+Our method outperforms state-of-the-art approaches in rendering quality across all 8 sequences, as well as in the overall average, surpassing both semantic and non-semantic SLAM baselines.
+The accurate semantic understanding learned by our system enhances both geometric and appearance representations through joint optimization, further improving the overall rendering fidelity of the SLAM system.
+The visual comparison of rendered images is shown in Fig.
+4.
+Benefiting from accurate semantic understanding and an improved SLAM framework, our method achieves significantly better detail preservation.
+Notable improvements include sharper and more accurate lamp boundaries in the first and fifth columns, as well as fine structural details, such as the shutters in the second column.
+Furthermore, our method produces high-quality renderings with fewer noise artifacts, evident in the clearer floor in the fourth column and the cleaner reconstruction of the chair in the sixth column.
+These results demonstrate that improved mapping quality directly contributes to enhanced rendering fidelity.
+The joint optimization of semantic and photometric losses proves mutually beneficial: semantic learning facilitates more accurate global 3D reconstruction, which in turn leads to better rendering performance in our Hier-SLAM++ system.
+Figure 4: Visualization of RGB image rendering performances on the Replica dataset [78].
+The first to third rows show the rendering outputs of the comparison methods.
+The fourth row presents the image rendering results of our Hier-SLAM++.
+The last row displays the ground truth images for reference.
+We use red boxes to highlight key differences regions within the image.
+The visual comparison clearly demonstrates that our method achieves superior rendering quality, closely resembling the ground truth images compared to state-of-the-art approaches.
+TABLE V: Rendering performance PSNR, SSIM, LPIPS on Replica. Best results are highlighted as first, second, third.
 Methods Metrics Avg. room0 room1 room2 office0 office1 office2 office3 office4
@@
 LPIPS $\downarrow$ 0.066 0.064 0.061 0.049 0.056 0.043 0.076 0.081 0.098
+IV-E Running time
+The runtime results for the compared methods are provided in Tab.
+VI, covering both NeRF-based and Gaussian Splatting-based SLAM methods.
+For the reason that one compared baseline using a flat representation (Hier-SLAM++-flat) cannot run successfully on an NVIDIA RTX 4090 due to GPU memory limitations, we report the runtime separately for RTX 4090 and L40S.
+In the first block of Tab.
+VI, we observe that, compared to NeRF-based methods [59], Gaussian-based state-of-the-art methods [16] achieve significantly higher operating speeds, benefiting from the fast rendering and efficient optimization capabilities inherent to Gaussian Splatting representation.
+Furthermore, with the carefully refined forward and backward gaussian module introduced in our system, which serves as the core component of gaussian splatting-based SLAM systems, our method (Hier-SLAM++- w/o sem) achieves up to 2.4× faster tracking and 2.2× faster mapping compared to the best-performing method [16].
+In the second block of Tab.
+VI, we evaluate the impact of different semantic representations on runtime performance.
+Our method maintains high efficiency by employing hierarchical semantic coding (Hier-SLAM++-one-hot), achieving nearly 3× faster tracking compared to the flat semantic representation (Hier-SLAM++-flat).
+Furthermore, our proposed binary semantic representation (Hier-SLAM++-binary) demonstrates even greater efficiency, enabling more lightweight performance.
+The flat version is slightly faster in mapping, with an improvement of 0.38 seconds per-frame, this is due to its cross-entropy loss calculation directly without hierarchical optimization for the proposed methods.
+Notably, our Hier-SLAM++ achieves a rendering speed of 2000 FPS, which increases to 3000 FPS when semantic information is not included.
 TABLE VI: Runtime on Replica/R0. Best results are highlighted as first.
@@
 Hier-SLAM++ (one-hot) 37.63 194.78 1.50 11.69
+Hier-SLAM++ (binary) 31.22 92.63 1.25 5.56
 L40S Hier-SLAM [25] 61.23 170.30 2.45 10.22
@@
   • The units are as follows: Tracking/Iteration (ms), Mapping/Iteration (ms), Tracking/Frame (s), and Mapping/Frame (s).
-TABLE VII: Semantic performance mIoU (%) and Parameter usage (MB) on Replica. Evaluations are conducted over a total of 102 semantic classes. Results are highlighted as first, second.
-Methods Avg. R0 R1 R2 Of0
-mIoU (%) NIDS-SLAM [12] 82.37 82.45 84.08 76.99 85.94
-DNS-SLAM [7] 84.77 88.32 84.90 81.20 84.66
-Hier-SLAM [25] 76.44 76.62 78.31 80.39 70.43
-Hier-SLAM++ (flat) 90.35 91.21 90.62 89.11 90.45
-Hier-SLAM++ (one-hot) 89.41 86.38 89.26 91.55 90.46
-Hier-SLAM++ (binary) 81.27 82.77 73.87 89.64 78.80
-SGS-SLAM [23] 92.72 92.95 92.91 92.10 92.90
-SemGauss-SLAM [24] 96.34 96.30 95.82 96.51 96.72
-SNI-SLAM [11] 87.41 88.42 87.43 86.16 87.63
-Param (MB) Hier-SLAM [25] 910.50 793 1126 843 880
-Hier-SLAM++ (flat) 2662.25 2355 3072 2560 2662
-Hier-SLAM++ (one-hot) 927.50 814 1126 867 903
-Hier-SLAM++ (binary) 591.75 528 690 563 586
-  • The gray background indicates that the different evaluation methods are used, where evaluations are conducted on a subset of the total classes.
 IV-F Hierarchical semantic understanding
-We perform semantic understanding experiments on the synthetic dataset Replica [41] to demonstrate the comprehensive performance of our proposed method.
-Replica [41] is a synthetic indoor dataset that includes 102 semantic classes with high-quality semantic annotations.
-Similar to previous works [24, 7, 11], we evaluate the mIoU performance as the quantitative semantic results in Tab.
-VII, where mIoU (%) is reported by comparing the rendered semantic map with the ground truth across all original semantic classes (102 classes).
-To further analyze the storage requirements of semantic SLAM methods, we report the parameter storage usage in the second block of Tab.
-VII.
+[…]
 From Tab.
@@
 The training time shown in Tab.
-VI demonstrates that our hierarchical version (Hier-SLAM++-one-hot) requires only 37% of the tracking time per frame while maintaining similar frame mapping performance compared to the flat version (Hier-SLAM++-flat).
+VI demonstrates that our hierarchical version (Hier-SLAM++ (one-hot)) requires only 37% of the tracking time per frame while maintaining similar frame mapping performance compared to the flat version (Hier-SLAM++ (flat)).
 The comparable mIoU performance, along with significantly reduced storage and computational cost, clearly demonstrates the effectiveness and efficiency of our proposed method.
-Furthermore, our binary version (Hier-SLAM++-binary) achieves state-of-the-art semantic understanding while achieves the best parameter storage efficiency, reducing storage usage to 63% of the one-hot version and requiring only 20% of the storage compared to the flat version.
+Furthermore, our binary version (Hier-SLAM++ (binary)) achieves state-of-the-art semantic understanding while achieves the best parameter storage efficiency, reducing storage usage to 63% of the one-hot version and requiring only 20% of the storage compared to the flat version.
 This reduction in storage usage highlights the efficiency of our proposed method.
-Because SGS-SLAM [23], SemGauss-SLAM [24] and SNI-SLAM [11] use different evaluation methods that restrict the mIoU calculation to a subset of the global semantic classes.
-Specifically, SGS-SLAM computes the average IoU for the semantic classes visible in each frame, and SemGauss-SLAM together with SNI-SLAM [11] evaluates performance based on 52 visible classes across all evaluation sequences, making direct comparisons unfair.
+Since works [11, 68, 67] report mIoU only on a subset of classes, making direct comparison unfair, we follow the same evaluation protocol as [67], which computes mIoU using only the visible semantic classes in each frame.
+Our method achieves an mIoU of $96.79\%$ with a storage usage of $910.50$ MB, demonstrating on-par semantic rendering performance with state-of-the-art methods [68], while maintaining efficient storage usage.
+Meanwhile, given that [68] benefits significantly from a large foundation model pre-trained on much larger and more diverse datasets, our results demonstrate that comparable performance can be achieved without relying on such heavy models.
 […]
@@
 The last row visualizes the semantic ground truth for comparison.
+TABLE VII: Semantic performance mIoU (%) and Parameter usage (MB) on Replica. Results are highlighted as first, second.
+Methods Avg. R0 R1 R2 Of0
+mIoU (%) total 102 classes NIDS-SLAM [12] 82.37 82.45 84.08 76.99 85.94
+DNS-SLAM [7] 84.77 88.32 84.90 81.20 84.66
+Hier-SLAM [25] 76.44 76.62 78.31 80.39 70.43
+Hier-SLAM++ (flat) 90.35 91.21 90.62 89.11 90.45
+Hier-SLAM++ (one-hot) 89.41 86.38 89.26 91.55 90.46
+Hier-SLAM++ (binary) 81.27 82.77 73.87 89.64 78.80
+mIoU (%) subset classes SNI-SLAM [11] 87.41 88.42 87.43 86.16 87.63
+SemGauss-SLAM [68] 96.34 96.30 95.82 96.51 96.72
+SGS-SLAM [67] 92.72 92.95 92.91 92.10 92.90
+Hier-SLAM++ (one-hot) 96.79 96.93 96.78 96.92 96.53
+Hier-SLAM++ (binary) 95.26 96.21 92.62 96.34 95.85
+Param (MB) Hier-SLAM [25] 910.50 793 1126 843 880
+Hier-SLAM++ (flat) 2662.25 2355 3072 2560 2662
+Hier-SLAM++ (one-hot) 927.50 814 1126 867 903
+Hier-SLAM++ (binary) 591.75 528 690 563 586
+  • Ours represents our method with a hierarchical representation, evaluated on a subset of semantic classes, consistent with [67].
+TABLE VIII: Monocular quantitative results on the TUM RGB-D dataset. Best results are highlighted as first, second.
+Sequences fr1-desk fr2-xyz
+Loop Method RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+w/ DROID-SLAM [84] 1.80 - - - 0.50 - - -
+ORB-SLAM2 [29] 2.00 - - - 0.60 - - -
+ORB-SLAM3 [30] 1.89 - - - 0.29 - - -
+Photo-SLAM [19] 1.54 20.97 0.743 0.228 0.98 21.07 0.726 0.166
+[…]
+MonoGS [17] 4.15 21.02 0.703 0.362 4.79 22.21 0.728 0.288
+Hier-SLAM++ 6.20 21.91 0.836 0.269 3.22 24.67 0.946 0.110
+Figure 6: Visualization of the established semantic 3D map across multiple levels, illustrating a coarse-to-fine semantic understanding of the complex scene.
+The bottom of the figure presents localization, rendering, and depth performance, offering a comprehensive overview of Hier-SLAM++’s effectiveness and demonstrating the scaling-up capability of our proposed method.
 IV-G Scaling up capability
@@
 As visualized in Fig.
-5, our estimated 3D global semantic map at different levels demonstrates a coarse-to-fine semantic understanding, showcasing our method’s scaling-up capability in handling real-world complex scenes.
-TABLE VIII: Monocular quantitative results on the TUM RGB-D dataset. Best results are highlighted as first, second.
-Sequences fr1-desk fr2-xyz
-Loop Method RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-w/ DROID-SLAM [48] 1.80 - - - 0.50 - - -
-ORB-SLAM2 [29] 2.00 - - - 0.60 - - -
-Photo-SLAM [19] 1.54 20.97 0.743 0.228 0.98 21.07 0.726 0.166
-[…]
-MonoGS [17] 4.15 21.02 0.703 0.362 4.79 22.21 0.728 0.288
-Hier-SLAM++ 6.20 21.91 0.836 0.269 3.22 24.67 0.946 0.110
+6, our estimated 3D global semantic map at different levels demonstrates a coarse-to-fine semantic understanding, showcasing our method’s scaling-up capability in handling real-world complex scenes.
 IV-H Monocular performance
@@
 VIII.
-For camera tracking, our method achieves competitive performance compared to state-of-the-art approaches without loop closure.
-Specifically, compared to the state-of-the-art Gaussian Splatting-based methods [17], our method exhibits similar tracking performance while achieving superior rendering quality.
-For SLAM methods with loop closure, which typically demonstrate superior camera tracking performance due to well-executed loop closure mechanisms that significantly enhance tracking accuracy and overall SLAM performance, our method still achieves superior rendering accuracy compared to Photo-SLAM [19], which incorporates loop closure.
-This highlights the strong rendering capabilities of our proposed method.
-With the assistance of the geometric prior provided by the feed-forward method, the joint optimization of both photometric loss and aligned depth loss contributes to improved rendering performance in the whole SLAM system.
-Figure 5: Visualization of the established semantic 3D map across multiple levels, illustrating a coarse-to-fine semantic understanding of the complex scene.
-The bottom of the figure presents localization, rendering, and depth performance, offering a comprehensive overview of Hier-SLAM++’s effectiveness and demonstrating the scaling-up capability of our proposed method.
+For SLAM methods without loop closure, our approach achieves competitive results compared to state-of-the-art methods in the same setting.
+For methods with loop closure, where well-executed loop mechanisms typically boost tracking accuracy and overall SLAM performance, our method still delivers better rendering accuracy compared to Gaussian-based SLAM systems [19].
+With the assistance of the geometric prior provided by the feed-forward model, the joint optimization of both photometric loss and aligned depth loss contributes to improved rendering performance across the entire SLAM system.
 IV-I Ablation study
-We evaluate SLAM performance under different tree designs in Tab.
+In this section, we evaluate the SLAM performance under different hierarchical tree designs.
+The results are presented in Tab.
 IX.
-The evaluation is conducted using the baseline method, Hier-SLAM [25], ensuring that any performance differences are solely attributable to the tree structures.
-In Tab.
-IX, each metric represents the average result from two runs on the first four sequences—Room0, Room1, Room2, and Office0—from the Replica dataset.
-We generate two different 5-level trees using our proposed tree generation procedure for the Replica dataset and incorporate them into the semantic SLAM system.
-As shown in Tab.
-IX, all methods demonstrate good performance, which can be attributed to the well-designed tree structures generated through both LLMs and 3D Generative Models.
-Additionally, the fully optimized hierarchical semantic loss contributes to the robustness of SLAM performance.
-Among all compared methods, tree-1 achieves the best semantic mIoU performance, the highest depth mapping accuracy, and close rendering quality (PSNR, SSIM, and LPIPS) compared with other settings.
-Therefore, we select tree-1 as the hierarchical tree structure leveraged within our semantic SLAM system.
-TABLE IX: Performance of different trees on Replica. Best results are highlighted in bold.
+Specifically, the experiments are conducted using our baseline setup, which adopts a one-hot hierarchical semantic representation and a naïve decoder with a single Softmax layer, denoted as Baseline in the table.
+This setup ensures that any observed performance differences are solely attributable to the tree structure itself.
+Each metric in Tab.
+IX represents the average result over two runs on the first four sequences—Room0, Room1, Room2, and Office0—from the Replica dataset.
+We generate two distinct 5-level hierarchical trees (tree1 and tree2) using our proposed tree construction framework, which integrates both LLM-generated semantics and 3D generative priors, and evaluate them within our semantic SLAM system.
+For comparison, we also include our previous Hier-SLAM [25] (tree0), which relies solely on LLM-generated semantics.
+As shown in the table, all configurations yield strong performance, which we attribute to the quality of the tree structures produced by the joint LLM and 3D prior-based generation process.
+In addition, the fully optimized hierarchical semantic loss contributes to the robustness of SLAM performance across tree variants.
+Among the compared configurations, tree2 achieves the highest semantic mIoU, the best depth reconstruction accuracy, and comparable rendering quality (in PSNR, SSIM, and LPIPS) relative to the alternatives.
+Therefore, we adopt tree2 as the default hierarchical structure in our semantic SLAM framework.
+TABLE IX: Performance of different trees on Replica. BEST RESULTS ARE HIGHLIGHTED AS FIRST.
 […]
-Hier-SLAM [25] 75.517 0.292 0.416 35.885 0.982 0.057
-[…]
-Hier-SLAM++${}^{\text{tree2}}$ 75.061 0.280 0.413 35.792 0.982 0.057
+Baseline (tree0) 75.517 0.292 0.416 35.885 0.982 0.057
+Baseline (tree1) 75.061 0.280 0.413 35.792 0.982 0.057
+Baseline (tree2) 76.754 0.302 0.408 35.825 0.982 0.058
+We also evaluate our proposed one-hot and binary semantic embeddings in previous experiments, including camera tracking performance in Tab.
+I and Tab.
+II, reconstruction accuracy in Tab.
+IV, rendering quality in Tab.
+V, as well as runtime and semantic understanding in Tab.
+VI and Tab.
+VII, respectively.
+For camera tracking, mapping, and rendering performance, both embeddings exhibit similar results.
+However, for runtime and semantic understanding, the two embeddings show differences.
+Specifically, our one-hot representation achieves better semantic understanding (higher mIoU in Tab.
+VII), while the binary embedding shows an 8% drop in mIoU on the full semantic class evaluation and less than 1% drop on the subset evaluation, but provides faster operation and significantly reduced storage—approximately 36% savings, as illustrated in Tab.
+VI and Tab.
+VII.
+As introduced in Section III, our proposed semantic embeddings provide two types of compactness: the one-hot representation effectively compresses semantic information, achieving faster training and lower memory usage, while the binary embedding further reduces storage and computation time at the cost of a small decrease in semantic accuracy (which remains minimal on subset evaluations).
+For tracking, mapping, and rendering, the performance difference between the two embeddings is negligible, as both integrate semantic information into the SLAM process with interpretable compactness derived from our carefully designed hierarchical tree representation.
 V Conclusion
@@
 Specifically, we proposed a compact and generalized hierarchical representation that encodes both semantic meaning and geometric attributes (i.e., size and shape), leveraging the capabilities of LLMs and 3D generative models.
-To enhance global semantic understanding, we introduced an improved hierarchical semantic loss.
+To enhance global semantic understanding, we introduced a novel hierarchical semantic loss.
 Furthermore, we extended the system from an RGB-D-only SLAM to support monocular input, utilizing geometric priors derived from a feed-forward SLAM approach.
@@
 These advances establish Hier-SLAM++ as a robust solution for semantic 3D understanding across diverse environments.
-Figure 6: Visualization of RGB image rendering performances on the Replica dataset [41].
-The first row presents the image rendering results of our Hier-SLAM++.
-The second to fourth rows show the rendering outputs of the comparison methods.
-The last row displays the ground truth images for reference.
-We use red boxes to highlight key differences regions within the image.
-The visual comparison clearly demonstrates that our method achieves superior rendering quality, closely resembling the ground truth images compared to state-of-the-art approaches.
 APPENDIX
-I Appendix 1. Implementation Details for Tree Generation
+I Implementation Details for Tree Generation
 During the hierarchical tree generation process, we utilize LLMs to group flat semantic classes at different levels based on physical size, semantic function, and to summarize and refine geometric grouping, ensuring balanced, meaningful, and descriptive outputs.
@@
 Always prefer singular shapes names like box, rectangle, soft, flat, etc.
-No Dublications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
+No Duplications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
 The group names must strictly by shapes and DO NOT leave any group empty, you could remove it if its empty.
@@
 New Group: If it is necessary to create a new group, feel free to do so, DO NOT create any non-original items.
-No Dublications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
+No Duplications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
 The output must be the same JSON format as below:
@@
     "<GROUP_2>": ["<ITEM_2>", ...],
+II Generated tree demonstration
+We demonstrate the generated hierarchical tree for the Replica dataset [78] in Fig.
+1.
+Based on the semantic categories in Replica, we leverage the capabilities of large language models and 3D generative models to construct a hierarchical tree.
+This tree represents semantic information as symbolic nodes (illustrated as colored nodes), which are used to produce compact semantic embeddings for each Gaussian primitive.
+These embeddings are then learned end-to-end across multiple views during SLAM operation.
+Figure 1: Visualization of the generated hierarchical tree on the Replica dataset [78].
+Different colors represent different tree depths.
+Each node corresponds to a symbolic concept, with label names shown next to the nodes.